### v1

# Field Matters: A lightweight LLM-enhanced Method for CTR Prediction

Abstract
Click-through rate (CTR) prediction is a fundamental task in modern recommender systems. In recent years, the integration of large language models (LLMs) has been shown to effectively enhance the performance of traditional CTR methods. However, existing LLM-enhanced methods often require extensive processing of detailed textual descriptions for large-scale instances or user/item entities, leading to substantial computational overhead. To address this challenge, this work introduces LLaCTR, a novel and lightweight LLM-enhanced CTR method that employs a field-level enhancement paradigm. Specifically, LLaCTR first utilizes LLMs to distill crucial and lightweight semantic knowledge from small-scale feature fields through self-supervised field-feature fine-tuning. Subsequently, it leverages this field-level semantic knowledge to enhance both feature representation and feature interactions. In our experiments, we integrate LLaCTR with six representative CTR models across four datasets, demonstrating its superior performance in terms of both effectiveness and efficiency compared to existing LLM-enhanced methods. Our code is available at https://anonymous.4open.science/r/LLaCTR-EC46.
1 Introduction
Click-through rate (CTR) prediction [1] is a core task in modern recommender systems (RS), with the objective of estimating the likelihood that a user will click on a specific item by modeling complex feature interactions. With the advent of Large Language Models (LLMs) and their remarkable abilities in content comprehension and semantic reasoning [2, 3], there has been a surge of interest in leveraging LLMs for CTR prediction. Recent studies mainly explored two paradigms: 1) LLMs as CTR predictors [4, 5, 6], where LLMs are either prompted or fine-tuned to directly perform CTR prediction; 2) LLM-enhanced CTR models [7, 8, 9, 10], where LLMs are leveraged to augment traditional CTR models by injecting semantic knowledge.
Despite their encouraging performance, these approaches encounter significant practical limitations, particularly regarding computational efficiency and economic feasibility. The high computational demands of LLMs make their direct deployment for online CTR prediction largely impractical, as they introduce considerable inference latency that violates real-time serving requirements. Although the second paradigm alleviates inference latency by retaining conventional CTR models for online serving, it still suffers prohibitive training costs. Specifically, these methods typically operate at the instance level [7, 9, 11] or the user/item level [8, 12, 10], requiring LLMs to process detailed textual descriptions for large-scale data instances or user/item entities. Empirically, we find that existing LLM-enhanced CTR methods (e.g., KAR[8], LLM-CF[9], CTRL[7] and EASE [10]) require over 290 times (on average) more computation time than baseline models on typical Amazon Video Games and MovieLens-1M datasets with millions of interactions (Figure 4 and Appendix B.1). This computational burden becomes even more pronounced when scaling to industrial settings involving billions of instances. These challenges naturally motivate a critical research question: How can we leverage LLMs to enhance CTR models in a more efficient and cost-effective manner?
To address this challenge, we explore a novel field-level enhancement paradigm that utilizes LLMs to extract semantic knowledge from feature fields to enhance CTR models. As illustrated in Figure 1, fields denote categories of features in CTR prediction. This paradigm offers two principal advantages: 1) Computational Efficiency: The number of fields is orders of magnitude lower than that of instances or entities (e.g., thousands vs. billions in industrial RS), yielding a drastic reduction in LLM computational overhead. 2) Crucial Semantic Information: Conventional CTR methods primarily utilize field IDs as categorical indicators, often neglecting the rich semantic information inherent in field descriptions. In fact, field semantics can significantly enhance both feature representation and feature interaction modeling — the two core components contributing the success of CTR prediction. For example, recognizing that the feature “4.7” belongs to the field “average rating” provides insight into item quality; while understanding the relationship between fields such as “user income” and “item price” can inform the importance of feature interactions. Although field-level semantic knowledge may not be as rich as cumbersome instance-level semantic information, its conciseness and importance offer substantial potential for improving model performance.
Motivated by these analyses, we introduce LLaCTR, a novel Lightweight LLM-enhanced CTR method through field-level enhancement. LLaCTR comprises two key components:
  • Self-supervised Field-feature Fine-tuning. While LLMs are pre-trained on open-domain corpora, they may lack domain-specific knowledge relevant to recommender systems. This limitation can hinder the quality of the distilled field semantic knowledge. To address this, we design a self-supervised task where the LLM is prompted to predict the field to which a given feature belongs. This fine-tuning leverages rich domain knowledge encoded in feature-field relationships, enabling the LLM to better understand the field semantics. Notably, since the number of fields is small and fine-tuning only requires a limited sample of features, this approach is more efficient than directly extracting instance or user/item-level knowledge as in prior work.
  • Field Semantic-guided Enhancement. We further utilize the field semantic knowledge (i.e., field semantic embeddings) distilled from LLMs to enhance traditional CTR models in two key aspects: 1) field embeddings are utilized to guide the learning of feature embeddings via a specific alignment loss, injecting semantic knowledge into feature representations; 2) field embeddings are transformed into field interaction matrix through a dedicated network to supplement and enhance feature interaction modeling.
Notably, LLaCTR is both flexible and lightweight, allowing for seamless integration into existing CTR models as a plug-and-play enhancement. In our experiments, we incorporate LLaCTR into six conventional CTR methods and observe significant performance improvements across four real-world datasets (with an average increase of 2.24%). Moreover, LLaCTR outperforms recent LLM-enhanced methods while incurring substantially lower computational overhead (by a factor of 10–100 times).
Figure 1: Current LLM-enhanced CTR paradigm versus our Field-level enhancement paradigm.
2 Preliminaries
Task Formulation. This work focuses on click-through rate (CTR) prediction [13, 14], a core task in many personalized services (e.g., recommender systems). Let $\mathcal{D}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N}$ denote the training dataset of user-item historical interaction records, where $\mathbf{X}_{i}=[\mathbf{x}_{i1},\mathbf{x}_{i2},\ldots,\mathbf{x}_{iK}]$ represents the input features across $K$-field for the $i$-th instance, and $\mathbf{x}_{ik}$ denotes the feature of the $k$-th field for the instance. In general, $\mathbf{x}_{ik}$ is a vector that can represent categorical, multi-valued, numerical features, etc. The label $y_{i}\in\{1,0\}$ indicates whether the user has clicked on the item in this instance. The goal of CTR is to learn a model from $\mathcal{D}$ that can accurately predict the click probability of a new instance.
Traditional CTR Models. Traditional CTR models [15, 16, 17, 18, 19, 20] primarily adopt deep learning paradigms and can generally be abstracted into three core components:
1) Feature Embedding Layer. Given the high-dimensional and sparse nature of raw input features, feature embedding is commonly employed to map the raw features into dense representations:
$\mathbf{E}_{i}=\text{EmbeddingLayer}(\mathbf{X}_{i}),$ (1)
where $\mathbf{E}_{i}=[\mathbf{e}_{i1},\mathbf{e}_{i2},\ldots,\mathbf{e}_{iK}]\in\mathbb{R}^{K\times D}$ represents the learned feature embeddings of the instance $i$, with $\mathbf{e}_{ik}$ denoting the D-dimensional embedding vector for $k$-th field feature.
2) Feature Interaction Layer. This crucial component is designed to effectively capture the complex feature interactions:
$\Phi(\mathbf{X}_{i})=\text{FeatureInteraction}(\mathbf{X}_{i},\mathbf{E}_{i}),$ (2)
where $\Phi(\cdot)$ is the hidden representation learned from feature interactions. Among various architectures, the most representative one is the factorization machine (FM) [15], which explicitly model bi-level feature interactions:
$\Phi_{FM}(\mathbf{X}_{i})={w_{0}}+\sum\limits_{k=1}^{K}{{\mathbf{w}_{k}}^{T}{\mathbf{x}_{ik}}}+\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}\text{>}}}})},$ (3)
where $w_{0},\mathbf{w}_{k}$ denotes the learnable weights, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the inner product of two vectors; $\varsigma(.)$ denotes the sum of all elements in the matrix. For convenience, here we adopt vector notation to represent the formula of Factorization Machines (FM), which is mathematically equivalent to the original formulation presented in [15]. The features are explicitly bi-level interacted with $\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}$ and its contribution is controlled by their embeddings $\text{<}\mathbf{e}_{ik},\mathbf{e}_{il}\text{>}$.
Building on FM, recent works have explored various extensions [21, 22, 23], such as introducing field-aware matrices [24, 17, 18] or incorporating neural layers [16, 25]. Meanwhile, other studies have proposed novel architectures, such as convolution operator [26, 27] and self-attention [28, 29], to better capture complex feature interactions.
3) Prediction Layer. Finally, the prediction layer transforms $\Phi(\mathbf{X}_{i})$ into the model prediction $\hat{y_{i}}$:
$\hat{y_{i}}=\sigma[\text{PredictionLayer}(\Phi(\mathbf{X}_{i}))],$ (4)
where $\sigma(\cdot)$ is the sigmoid function. The model is typically optimized with the binary cross-entropy (BCE) loss [30] with:
$\mathcal{L}_{BCE}(y,\hat{y})=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\left(\hat{y_{i}}\right)+(1-y_{i})\log\left(1-\hat{y_{i}}\right)\right],$ (5)
While these traditional CTR models have made significant progress in the past few decades, they are often plagued by the issue of semantic information loss. Specifically, Some methods have utilized field IDs as categorical indicators [24, 17, 18], but the rich semantic information inherent in field descriptions has not been exploited. This limitation motivates us to leverage LLMs to enrich the traditional CTR models with field-level knowledge.
LLM-enhanced CTR Models. To fully exploit semantic knowledge, Large Language Models (LLMs) [2, 3] have been extensively investigated for enhancing traditional CTR models [8, 7, 9, 10, 31]. As illustrated in Figure 1, existing LLM-enhanced CTR methods primarily operate at the instance level or user/item level. Typically, these approaches first organize the features of users, items, or instances into textual descriptions, which are then provided as prompts to the LLM for reasoning or summarization. The resulting semantic knowledge, often at the user/item or instance level, is further encoded into semantic embeddings to augment traditional CTR models (see Section 5 for details). However, due to the large scale of users/items and instances in real-world applications, such strategies incur prohibitive computational costs during training, inference and encoding. For example, on representative datasets such as Amazon Video Games and MovieLens-1M, the total training time of existing LLM-enhanced CTR methods (e.g., KAR [8], LLM-CF [9], CTRL [7] and EASE [10]) is over 290 times (on average) that of standard CTR models on average (see Appendix B.1 for details). Moreover, the complex and voluminous semantic knowledge generated at the user/item or instance level is often difficult for traditional CTR models to effectively assimilate and utilize. Given these limitations, it is imperative to explore new paradigms for integrating semantic knowledge into CTR models.
3 Methodology
Figure 2: The overall framework of proposed LLaCTR. It contains 1) SSFT: leveraging self-supervised field-feature fine-tuning to improve LLMs’ ability to capture field semantics; 2) FRE: leveraging field semantic embeddings to guide the learning of the feature representation; 3) FIE: leveraging field semantic embeddings to enhance the models of feature interactions.
In this section, we introduce LLaCTR, which enhances CTR prediction through field-level semantic knowledge from LLMs. As shown in Figure 2, LLaCTR first employs self-supervised field-feature fine-tuning to improve LLMs’ ability to capture field semantics (Section 3.1), and then leverage the field knowledge distilled from LLMs to enhance both feature representation (Section 3.2) and feature interactions (Section 3.3).
3.1 Self-supervised Field-feature Fine-tuning (SSFT)
LLMs are pre-trained on open-domain natural language corpora [2, 3], which often lack domain-specific knowledge relevant to CTR prediction. This limitation can hinder their ability to distill high-quality field semantic knowledge. Specifically, LLMs may not fully understand the meaning of field descriptions in CTR tasks or their relationships with features. To address this, we propose a self-supervised fine-tuning strategy that prompts LLMs to predict the field to which a given feature belongs, thereby enhancing their understanding of field semantic.
Specifically, let $F_{k}$ represent the descriptions of the $k$-th field, and $f^{k}_{j}$ denote the description of the $j$-th feature within field $k$. As shown in Figure 3, we construct self-supervised training instances $(\mathcal{P}(f^{k}_{j}),F_{k})$, where $\mathcal{P}(f^{k}_{j})$ is a prompt containing the task description, feature descriptions, and candidate field descriptions, and $F_{k}$ denotes the $k$-th field description as the target response. These prompt-response pairs are used to fine-tune LLMs with language generative loss:
$\mathcal{L}_{LG}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log P_{LLM}\left(F_{k}\mid\mathcal{P}(f^{k}_{j})\right),$ (6)
where $P_{LLM}\left(F_{k}\mid\mathcal{P}(f^{k}_{j})\right)$ represents the LLM’s probability of generating the correct field description $F_{k}$ given the prompt $\mathcal{P}(f^{k}_{j})$. This loss improves the LLM’s ability to capture field-feature correlations, injecting domain knowledge and enabling the generation of higher-quality semantic representations.
Besides the language generative loss, we find the contrastive loss [32] is also effective:
$\mathcal{L}_{CL}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log\frac{\exp(cos(En(\mathcal{P}(f_{j}^{k})),En(F_{k}))/\tau)}{\sum_{l=1}^{K}\exp(cos(En(\mathcal{P}(f_{j}^{k})),En(F_{l}))/\tau)},$ (7)
where $cos(\cdot)$ denotes the cosine similarity function, $\tau$ is the temperature parameter and $En(\cdot)$ are the LLM’s encodings of the language descriptions, respectively.
The contrastive loss is designed to align the semantic embeddings of prompts with their corresponding answers, thereby effectively injecting feature-field correlation knowledge in embeddings. Since the contrastive loss operates directly on embeddings, which are subsequently used to enhance CTR models, this work prioritizes the use of contrastive loss and empirically find slightly better performance than generative loss (Appendix B.2).
After fine-tuning, the LLM’s encodings of field descriptions are distilled:
$\textbf{h}_{k}=En(F_{k}).$ (8)
For convenience, we collect the field embeddings as a matrix $\textbf{H}=\{\textbf{h}_{k}\}_{k=1}^{K}$. These field embeddings will further utilized to enhance traditional CTR models.
Figure 3: The template of self-supervised tuning.
Notably, such field-level operations are highly efficient, as the LLMs require to process only a small number of fields (typically less than 100) rather than millions of instances or user/item pairs in the experimental datasets. This difference in scale is even more pronounced in real-world industrial recommender systems, where there may be thousands of fields, but hundreds of millions of users/items and billions of instances. Some readers may be concerned about the potential inefficiency of self-supervised fine-tuning. However, this process is also lightweight, as it only requires sampling a small number of features per field for training. For example, sampling just 500 features per field is sufficient to achieve significantly better performance compared to recent LLM-enhanced methods, while incurring substantially lower training time. Empirical comparisons of efficiency can be found in Figure 4. The detailed time complexity analyses of LLaCTR and other LLM-enhanced baselines are provided in Appendix B.1.
3.2 Feature Representation Enhancement (FRE)
The quality of feature representations is critical for the success of CTR models. However, existing methods often overlook the semantic knowledge contained in field descriptions. To address this, we transfer the field semantic knowledge distilled from LLMs to enhance feature embeddings.
Given the embedding space gap between LLMs and CTR models, we first map original field embeddings $h_{k}$ into the embedding space of CTR models using a learnable adaptor:
$\textbf{h}^{{}^{\prime}}_{k}=\text{Adaptor}(\textbf{h}_{k}),$ (9)
where $\text{Adaptor}(\cdot)$ can be implemented as a simple linear transformation. We then align feature embeddings with their corresponding field embeddings through normalized KL-divergence [33]:
$\mathbf{h}^{{}^{\prime}}_{k}\leftarrow\text{Softmax}(\mathbf{h}^{{}^{\prime}}_{k}),\ \mathbf{e}_{ik}\leftarrow\text{Softmax}(\mathbf{e}_{ik}),$ (10)
$\mathcal{L}_{KL}=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{D}_{\text{KL}}(\mathbf{e}_{ik},\mathbf{h}^{{}^{\prime}}_{k})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{e}_{ik}\cdot\ln({\mathbf{e}_{ik}}/{\mathbf{h}_{k}^{{}^{\prime}}}).$ (11)
Specifically, here we employ normalized KL-divergence, as it eliminates the influence of embedding magnitude and focuses on aligning the underlying distributions. We also experimented with other commonly used alignment losses, such as contrastive loss, but found that normalized KL-divergence consistently yields superior performance (Appendix B.2). By minimizing $\mathcal{L}_{KL}$, feature embeddings are able to absorb semantic knowledge from the field embeddings, resulting in higher-quality representations. Intuitively, the field semantic embeddings serve as prototypes, encouraging the feature embeddings close to these semantic centers during training.
The overall training objective combines semantic alignment with the original CTR objective:
$\mathcal{L}=\mathcal{L}_{BCE}+\lambda_{kl}\mathcal{L}_{KL},$ (12)
where $\lambda_{kl}$ serves as a hyperparameter balancing the contributions of the two components.
3.3 Feature Interaction Enhancement (FIE)
Field information has been shown beneficial to improve feature interaction modeling in CTR tasks [24, 15, 18]. However, these traditional method only utilize field ID information by employing implicit embedding layers [24] or learnable parameters matrices [15, 18], without taking the textual information of fields into account. Field semantic information provides new insights into feature interactions. For example, semantically, interactions between the features belonging to "user income" and "item price" are likely highly useful for click prediction.
To leverage this insight, we first compute the importance of feature interactions of two fields based on their semantic embeddings:
$m_{ij}^{F}=\text{FieldInteraction}(\mathbf{h}_{i},\mathbf{h}_{j}),$ (13)
[…]
The field-aware importance scores derived from semantic embeddings are then utilized to guide the learning of feature interactions. Specifically, we explicitly incorporate these scores into a bi-level feature interaction modeling, thereby enhancing the feature interaction layer of CTR methods:
$\Phi^{new}(\mathbf{X}_{i})=\Phi(\mathbf{X}_{i})+\lambda_{fm}\cdot\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}^{T}\text{>}m_{kl}^{{}^{\prime}}}}})},$ (14)
where the field-aware importance score $m_{kl}^{{}^{\prime}}$ modulates the strength of feature interactions. Here, we augment the original feature interaction layer by integrating an additional field-aware feature interaction component, with $\lambda_{fm}$ controlling the relative contribution of the two components.
Notably, our proposed feature interaction enhancement strategy can be seamlessly integrated into a wide range of existing CTR models. This generality motivates our approach of introducing a supplementary field-aware feature interaction component, rather than modifying the inherent feature interaction architecture of each model. In our experiments, we incorporate our LLaCTR module into six representative CTR methods, and observe substantial performance improvements across almost all cases. Detailed results are presented in Section 4, and the implementation details are provided in the Appendix A.4.
4 Experiments
4.1 Experimental Setup
Datasets. We adopt four widely-used public datasets in LLM-enhanced CTR task: Amazon Gift Cards, Amazon Video Games, Amazon Digital Music and MovieLens-1M. We closely refer to [11, 8] and partition each dataset into training, validation, and testing sets with ratios of 8:1:1 according to the timestamp. The details of the above datasets are summarized in Appendix A.1.
Metrics. Following previous work [8, 34, 35], we employed two widely-used metrics LogLoss (binary cross-entropy loss) and AUC (Area under the ROC curve) to evaluate performance. RelaImpr [36], the relative AUC improvement compared to the basic CTR model [28, 7] is also reported. The readers may refer to Appendix A.2 for more details.
Baselines. For comparisons, we selected KAR (RecSys’24) [8], LLM-CF (CIKM’24) [9], CTRL (TORS’23) [7] and EASE (CIKM’24) [10] as LLM-enhanced baselines. These baselines are closely related and representative LLM-enhanced CTR methods. For fair comparisons, unless otherwise specified, we consistently used the Llama3-8B [3] as the LLM. We also include two representative PLM-enhanced methods - CELA (arXiv’24) [12] and ClickPrompt (WWW’24) [11] for comparisons.
We integrated these baseline methods into the following representative traditional CTR models: 1) Classic FM-based methods: FM (ICDM’10) [15] and DeepFM (IJCAI’17) [16]; 2) field-based CTR methods: FwFM (WWW’18) [17] and FmFM (WWW’21) [18]; 3) The state-of-the-art CTR methods: FinalMLP (AAAI’23) [19] and WuKong (ICML’24) [20]. The readers may refer to Appendix A.4 for more details about these methods.
Hyperparameter settings. A grid search is utilized to find the optimal hyperparameters. For all compared methods, we closely refer to the configurations provided in their respective publications to ensure their optimal performance. For LLaCTR, $\lambda_{kl}$ and $\lambda_{fm}$ are tuned in {0.01, 0.05, 0.1, 0.3, 0.5, 0.7, 1}, and the feature sampling number in fine-tuning is set at 1000. The detailed hyperparameter settings are provided in Appendix A.5.
4.2 Experimental Results
In this section, we primarily focus on performance comparison, efficiency analysis, and ablation studies. Additional experimental results, including hyperparameter sensitivity analyses and case studies, are provided in Appendix B.
Table 1: Performance comparisons of LLaCTR with existing PLM/LLM-enhanced methods on six CTR backbone models. The results of PLM-enhanced methods are reported on their best backbones. The best result is bolded and the blue-colored zone indicates that LLaCTR is better than the basic CTR backbone. The line "Average RelaImpr" indicates the average RelaImpr of LLaCTR over six CTR backbones. The mark "*" indicates the improvement is statistically significant ($p$-value $<0.05$).
[…]
KAR 0.4551 0.6643 1.34% 0.5803 0.6705 1.93% 0.3410 0.7607 -0.21% 0.4073 0.8358 0.13%
LLM-CF 0.4580 0.6666 2.77% 0.5805 0.6682 0.58% 0.3351 0.7633 0.78% 0.4128 0.8337 -0.50%
CTRL 0.4595 0.6636 0.93% 0.5802 0.6708 2.13% 0.3360 0.7611 -0.04% 0.4128 0.8336 -0.53%
EASE 0.4534 0.6667 2.84% 0.5804 0.6701 1.69% 0.3359 0.7625 0.48% 0.4101 0.8357 0.10%
[…]
KAR 0.4753 0.6778 3.80% 0.5871 0.6720 1.65% 0.3439 0.7587 0.18% 0.4207 0.8335 -0.20%
LLM-CF 0.4723 0.6736 1.36% 0.5769 0.6730 2.28% 0.3441 0.7588 0.21% 0.4140 0.8353 0.34%
CTRL 0.4746 0.6744 1.83% 0.5757 0.6758 3.92% 0.3432 0.7576 -0.26% 0.4142 0.8352 0.31%
EASE 0.4748 0.6739 1.52% 0.5807 0.6738 2.74% 0.3360 0.7604 0.83% 0.4180 0.8349 0.22%
[…]
KAR 0.4661 0.6694 0.19% 0.5791 0.6741 3.42% 0.3432 0.7575 0.88% 0.4020 0.8360 -0.15%
LLM-CF 0.4672 0.6698 0.47% 0.5806 0.6702 1.13% 0.3397 0.7567 0.57% 0.4044 0.8371 0.18%
CTRL 0.4679 0.6699 0.50% 0.5776 0.6755 4.28% 0.3414 0.7561 0.33% 0.4004 0.8372 0.21%
EASE 0.4661 0.6703 0.74% 0.5795 0.6765 4.84% 0.3415 0.7571 0.71% 0.4030 0.8369 0.12%
[…]
KAR 0.4595 0.6754 0.42% 0.5767 0.6711 0.29% 0.3347 0.7640 0.43% 0.4062 0.8332 0.21%
LLM-CF 0.4652 0.6740 -0.43% 0.5877 0.6709 0.20% 0.3317 0.7627 -0.04% 0.4066 0.8329 0.11%
CTRL 0.4593 0.6751 0.21% 0.5826 0.6714 0.47% 0.3310 0.7650 0.82% 0.4059 0.8332 0.21%
EASE 0.4590 0.6753 0.33% 0.5750 0.6729 1.35% 0.3311 0.7637 0.33% 0.4056 0.8334 0.27%
[…]
KAR 0.4881 0.6726 0.35% 0.5820 0.6817 3.19% 0.3429 0.7546 -0.41% 0.4094 0.8290 -0.33%
LLM-CF 0.4781 0.6731 0.62% 0.5845 0.6785 1.36% 0.3369 0.7560 0.13% 0.4041 0.8299 -0.06%
CTRL 0.4719 0.6781 3.55% 0.5835 0.6790 1.65% 0.3399 0.7561 0.15% 0.4033 0.8328 0.82%
EASE 0.4773 0.6736 0.93% 0.5776 0.6796 1.97% 0.3414 0.7560 0.12% 0.4049 0.8309 0.24%
[…]
KAR 0.4794 0.6721 -0.43% 0.5927 0.6758 -0.53% 0.3429 0.7506 -0.31% 0.4099 0.8316 -0.21%
LLM-CF 0.4756 0.6752 1.36% 0.5805 0.6788 1.15% 0.3379 0.7560 1.83% 0.4084 0.8293 -0.90%
CTRL 0.4753 0.6769 2.37% 0.5792 0.6800 1.84% 0.3399 0.7561 1.85% 0.4029 0.8367 1.31%
EASE 0.4748 0.6758 1.71% 0.5783 0.6801 1.91% 0.3381 0.7556 1.67% 0.4031 0.8334 0.33%
[…]
Comparing with traditional CTR models. From the Table 1, we observe LLaCTR can be applied to various types of basic CTR models, and yield performance gains in most cases. The improvements brought by LLaCTR are impressive, achieving an average AUC relative improvement of 2.24% over all basic CTR models across four datasets. These results demonstrate the effectiveness of LLaCTR by injecting field semantic knowledge.
Comparing with PLM/LLM-enhanced CTR models. We observe that LLaCTR outperformed two representative PLM-enhanced CTR models (CELA and ClickPrompt) and four LLM-enhanced CTR models (KAR, LLM-CF, CTRL and EASE) in most cases (89%), which validates the effectiveness of our field-level enhancement paradigm. This result is highly surprising, as we just use the lightweight field knowledge rather than the cumbersome instance-level (or user/item-level) knowledge utilized by recent work. The reason may be : 1) field knowledge is indeed important and helpful for CTR prediction; 2) the lightweight field knowledge has been effectively exploited, enhancing both feature representation and feature interactions.
Efficiency Analysis. The AUC performance and the total training time of compared methods are shown in Figure 4. The performance of LLM-enhanced methods are reported on their best backbones. We observe: 1) Compared to three representative LLM-enhanced methods, LLaCTR can improve efficiency by 10-100x and achieve better performance. 2) With the increase of sampling number in the SSFT module, both the training time and AUC of LLaCTR will increase. This indicates increasing the number of samples used for self-supervised learning can yield higher-quality field semantic knowledge, but require more training time cost. 3) Even without fine-tuning, LLaCTR still achieves decent results. This suggests our LLaCTR could be applied in resource-constrained scenarios, enhancing existing CTR methods with incurring limited additional time cost.
Ablation Study. We further conduct the ablation study, where the Self-supervised Fine-tuning (SSFT) module, Feature Representation Enhancement (FRE) module or Feature Interaction Enhancement (FIE) module has been removed respectively. The results are presented in Table 2. As can be seen, both the three components are important — with few exceptions, removing SSFT (w/o FT), FRE (w/o FRE) or FIE ($\lambda_{fm}$) module would result in performance drops.
Figure 4: AUC and training time of compared methods. The “sam-num” represents the sampling feature number for each field of LLaCTR, and "w/o FT" represents fine-tuning has been removed.
Table 2: Ablation study on LLaCTR with DeepFM and WuKong backbones. The results on other backbones are presented in Appendix B.3.
[…]
w/o FT 0.4687 0.6759 0.5771 0.6732 0.3393 0.7683 0.4064 0.8352
[…]
5 Related Work
Traditional CTR Prediction. CTR Prediction serves as a core function module in personalized online services [37, 1, 38, 39, 16, 40]. The key idea of traditional CTR prediction models is to capture the feature interaction patterns, which indicates the combination relationships of multiple features. The feature interaction layer is generally implemented by a combination of various operations. Product operators include classical bi-level models [15, 41, 42] and higher-order [43, 44, 45, 46, 47] models. Convolutional operators [48, 27, 26] utilize CNN [49] and GCN [50] to capture the implicit high order interaction. Due to the success of attention mechanism, attention-based methods [51, 21, 34, 35, 28, 29] also achieved great success. In recent years, several field-aware CTR methods have been proposed [24, 17, 18], however, these methods relied solely on field IDs as categorical indicators without leveraging the rich semantic information inherent in field descriptions.
LLMs as CTR predictors. Initial efforts explored the usage of pre-trained language models as CTR predictors by reformulating recommendation tasks as natural language [4, 52, 53]. With the advent of LLMs and their remarkable abilities in content comprehension and semantic reasoning [2, 3], recent efforts have shifted to explore using larger-scale Language Models to generate CTR prediction results [54, 55, 56, 57, 58, 59, 60, 31], e.g., TALLRec [5] uses instruction tuning with user’s historical interactions, enabling LLMs to achieve decent CTR prediction capability in zero-shot and few-shot scenarios, while HLLM [61] utilizes two independent LLMs to model users and items separately, and jointly trains them with both generative and discriminative recommendation losses. However, the high inference latency and resource demands of LLMs poses significant obstacles to their deployment in real-world large-scale recommender systems. Although recent studies have proposed various acceleration strategies, e.g., BAHE [6] freezes the shallow layers of LLMs to encode users’ atomic interactions and pre-stores them offline to reduce the number of parameters involved in training and inference, while Rella [62] reduces inference overhead by shortening user interaction sequences through the elimination of items that are semantically irrelevant to the target item, the latency and resource consumption nonetheless remain substantial due to the LLMs’ large parameter counts and complex architectures.
LLM-enhanced CTR Prediction. To address the efficiency challenges inherent in LLMs-based CTR predictors, the LLM-enhanced CTR paradigm has been extensively investigated as an alternative approach. This paradigm primarily focus on leveraging the powerful semantic encoding and knowledge reasoning capabilities of LLMs to enhance traditional CTR models [63, 8, 9, 7, 10]. For example, KAR [8] leverages the LLM as a open-word knowledge base to enhance the profile of users and items. LLM-CF [9] uses LLMs to generate a chains of thought (CoT) for sampled instances, which is further retrieved to enhance CTR models. CTRL [7] utilizes the LLM as the feature encoder for instances and uses a fine-grained contrastive learning scheme to align the knowledge. EASE [10] trains a semantic adaptor to align item embeddings with LLM, and use it for encoding during inference. Other work also focuses on specific CTR scenarios, such as cross-domain[64, 65] and interpretability [66, 67]. Despite their promising performance, they involve enhancements at the instance-level or user/item-level, which results in great inefficiency issue in dealing with large-scale instances or entities in practice. Our approach is applied on the more efficient and cost-effective field-level, without incurring superabundant computational overhead.
There are also two related topics: 1) PLM-enhanced CTR methods: Early studies have leveraged pretrained language models (PLMs, like BERT [68] and RoBERTa[69]) to enhance CTR models via encoding textual features [70, 71, 72, 73, 53, 11, 12], e.g., CELA [12] pretrains a PLM on item descriptions and aligns item textual embeddings with ID-based embeddings for enhancement. ClickPrompt [11] uses a CTR model to encode instances into soft tokens, which are then treated as prefix tokens in a PLM for generating instance-level semantic knowledge. However, the capacity of early PLMs is generally inferior to that of recent LLMs, and recent studies have demonstrated the superior performance of LLM-enhanced CTR methods. Consequently, our work aims to further investigate and advance the state-of-the-art in LLM-enhanced CTR modeling. 2) LLM-enhanced recommendation: Recent years also witnessed the integration of LLMs on other recommendation tasks,e.g., sequential recommendation [74, 75, 59, 76, 77, 78], collaborative filtering [79, 80, 65, 81], cross-domain recommendation [82, 83]. It is important to note that CTR prediction differs from these tasks in that it typically involves various features and emphasizes modeling feature interactions. Thus, our LLaCTR is specifically designed to leverage LLMs for improved feature representation and interactions in CTR prediction.
6 Conclusion and Limitations
This work proposes a novel lightweight method, LLaCTR, which leverages LLMs to extract semantic knowledge from feature fields for enhancing CTR models. Specifically, LLaCTR employs a self-supervised field-feature fine-tuning to distill high-quality field-level knowledge, which is subsequently utilized to improve both feature representation and feature interaction modeling. Extensive experiments demonstrate the superiority of LLaCTR in both predictive accuracy and computational efficiency. Nevertheless, a current limitation of LLaCTR is that it leverages field-level semantic knowledge to guide bi-level feature interaction modeling. Exploring more advanced strategies for incorporating field knowledge to capture higher-order feature interactions remains an interesting and promising direction for future research.
Appendix A Experimental Details
A.1 Datasets
We conduct experiments on four real-world datasets from different recommendation scenarios. The datasets statistics are summarized in Table A.1.
Table A.1: Statistics of the datasets.
[…]
#Feature 1,981,330 60,119,995 1,695,642 9,001,881
#User 132,732 2,766,656 100,952 6,040
#Item 1,137 137,249 70,511 3,706
#Interaction 152,410 4,624,615 130,434 1,000,209
Density $1.01\times 10^{-3}$ $1.22\times 10^{-5}$ $1.83\times 10^{-5}$ $4.47\times 10^{-2}$bigstrut[b]
  • Amazon Review Datasets [^1] https://amazon-reviews-2023.github.io/ are widely-used e-commercial datasets [9, 7, 11, 84] from Amazon website with ratings ranging from 1 to 5. We use the Gift Cards, Video Games and Digital Music subsets in our experiments. Following the standard practice in Lin et al. [11], we binarize the ratings with a threshold of 4 and use 5-core setting, i.e., all users and items have at least 5 interactions.
  • MovieLens-1M Dataset [^2] https://grouplens.org/datasets/MovieLens/1m/ is a movie recommendation dataset from Movielens website with ratings ranging from 1 to 5. We binarize the ratings with a threshold of 4, while removing neutral samples with ratings equal to 3 following previous work [7].
Following [11], we organized the review behaviors in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1 after preprocessing. All the experiments are implemented with PyTorch and run on 4x Nvidia A800 GPUs.
A.2 Metrics
To evaluate the performance of CTR prediction methods, we closely follow [28, 8, 11, 7] and adopt AUC (area under the ROC curve), LogLoss (binary cross-entropy loss) and RelaImpr (AUC relative improvement) as the evaluation metrics:
  • AUC [85]: AUC is the area under the ROC curve and measures the goodness of order by ranking all the items with predicted CTR. For a random guesser, the value of AUC is 0.5.
  • LogLoss: Logloss is the logarithmic loss value of binary cross-entropy loss and quantifies the accuracy of predicted probabilities by penalizing false classifications.
  • RelaImpr [36]: RelaImpr measures the AUC relative improvement over basic CTR models and is defined as below:
    $RelaImpr=(\frac{\text{AUC (measured\ model)}-0.5}{\text{AUC (base\ model)}-0.5}-1)\times 100\%.$ (15)
Slightly higher AUC or lower Log Loss (e.g., 0.001) can be regarded as significant improvement in CTR prediction, as indicated by previous studies [34, 35].
A.3 Baselines
We reproduced the following LLM-enhanced CTR methods as baselines in our experiments:
  • KAR (RecSys’24) [8]: KAR leverages the LLM as a open-word knowledge base to enhance the profile of users and items, and then pass the encoded user/item embeddings through a knowledge adaptation network to enhance traditional CTR models.
  • LLM-CF (CIKM’24) [9]: LLM-CF uses LLMs to generate a base of chains of thought (CoT) for sampled instances, which are further retrieved to enhance CTR models. Consistent with the original paper, we sampled 10% of the instances as sampled data for fine-tuning and CoT data generation, and during downstream CTR prediction, we also retrieve the Top 4 samples closest in semantics from the CoT dataset as enhanced knowledge.
  • CTRL (TORS’23) [7]: CTRL utilizes the LLM as the feature encoder for instances and uses a fine-grained contrastive learning scheme to align cross-modal knowledge. In our implementation, we used Llama3-8B [3] as the LLM to encode instances embeddings for fair comparison.
  • EASE (CIKM’24) [10]: EASE trains a semantic adaptor network to align item embeddings with the LLM, and then use it for encoding during inference. Following the original paper, we inited the semantic adapter using BERT-Base [68] with 12 transformer blocks, and cross-attention layers are added to the 6-th and 12-th transformer blocks.
For fair comparisons, unless otherwise specified, we consistently used the Llama3-8B [3] as the LLM in all the LLM-enhanced CTR methods.
Besides, we also reproduced two representative PLM-enhanced methods:
  • CELA (arXiv’24) [12]: CELA first pretrains a PLM on item descriptions, then aligns item textual embeddings with pretrained ID-based embeddings, and finally merges them into CTR models for enhancement. We selected Robert-base [69] as the PLM as it performs the best in the original paper.
  • ClickPrompt (WWW’24) [11]: ClickPrompt uses a CTR model to encode instances into soft tokens, which are then treated as prefix tokens in a PLM for generating instance-level semantic knowledge. Following the original paper, we selected Robert-base [69] as the PLM in our experiment.
A.4 Backbones
We selected six representative CTR methods as backbones in our experiments, including:
  • FM (ICDM’10) [15]: FM introduces factorization machines to model pairwise feature interactions via factorized parameters.
  • DeepFM (IJCAI’17) [16]: DeepFM combines factorization machines with deep neural networks to jointly learn low-order and high-order feature interactions. We fix the DNN layers at [300, 300, 128] in our experiments.
  • FwFM (WWW’18) [17]: FwFM proposes field-weighted factorization machines that learn field-specific interaction weights to better capture heterogeneous feature relationships.
  • FmFM (WWW’21) [18]: FmFM enhances FM with field-matrix factorization using learnable projection matrices per field pair, enabling more expressive cross-field interactions.
  • FinalMLP (AAAI’23) [19]: FinalMLP designs two Stream Feature Interaction networks to progressively refine feature representations learned by MLP. We select the MLP hidden units size from {400, 500} and the MLP layer number from {2, 3}, respectively.
  • WuKong (ICML’24) [20]: WuKong employs an interaction stack on the WuKong layer to capture feature interactions, where each WuKong layer consists of a Factorization Machine Block (FMB) and a Linear Compress Block. We select the interaction layer number from {4, 8}, the FMB size from {128, 200}, the FMB layer from {2, 3} and the compression dimension from {24, 32} respectively.
For the implementation of LLaCTR, the SSFT module and FRE module remained consistent across all backbones. The implementation of the FIE module had slightly difference on explicit and implicit feature interaction backbones. Specifically:
  • 1) For explicit feature interaction backbones (FM, DeepFM, FwFM and FmFM), the FIE module was implemented by adding the learnable field interaction matrix on the original feature interaction matrix. By explicitly guiding the learning of feature interactions modeling, the FIE module can directly inject the field interaction knowledge into the CTR models.
  • 2) For implicit feature interaction backbones (FinalMLP and WuKong), the FIE module was treated as an additional plugin network to explicitly influence the second-order feature interactions learning and generate CTR prediction logit, which was later fused into the final prediction result of the original CTR model.
The hyperparameter settings of each method are detailed in Appendix A.5. For all the backbone models, we reused the implement of an open-source CTR prediction library [^3] https://github.com/reczoo/FuxiCTR - FuxiCTR [86], following the previous work [19, 87, 47].
A.5 Hyperparameters
A.5.1 Hyperparameter Settings
We use Adam [88] optimizer to train all the CTR models. The learning rate (lr) is set as 0.001 and the weight decay (wd) is tuned in {1e-2, 1e-3, 1e-4, 1e-5, 1e-6, 0}. We fix the feature embedding size to 32 and batch size to 4096 for all the backbones and datasets. For LLaCTR fine-tuning, we fix the batch size as 128, the learning rate of LLM as 1e-4 and rank of LoRA [89] as 8, respectively. We set the feature sampling number as 1000 by default, and set temperature parameter $\tau$ of contrastive loss into 0.02 following previous work [90]. During LLaCTR enhancement, the $\lambda_{kl}$ and $\lambda_{fm}$ are tuned in {0.01, 0.05, 0.1, 0.3, 0.5, 0.7, 1}.
A.5.2 Optimal Hyperparameters
The hyperparameters we searched include: the $\lambda_{kl}$ for Feature Interaction Enhancement (FRE) module, the $\lambda_{fm}$ for Feature Interaction Enhancement (FIE) module, the weight decay (wd) of Adam [88] optimizer and other hyperparameters for backbone CTR models. Table A.2 shows the optimal hyperparameters, including four datasets (Gift Cards, Video Games, Digital Music and MovieLens-1M) and six backbones (FM, DeepFM, FwFM, FmFM, FinalMLP and WuKong).
Table A.2: Optimal hyperparameters for 4 datasets and 6 backbones.
[…]
$\lambda_{fm}$ 0.01 0.01 0.7 0.5
[…]
$\lambda_{fm}$ 0.5 0.5 0.3 0.7
[…]
$\lambda_{fm}$ 0.1 0.1 0.3 0.3
[…]
$\lambda_{fm}$ 0.1 0.1 0.05 0.7
[…]
$\lambda_{fm}$ 0.1 0.3 0.01 0.05
wd 0.01 0.00001 0.000001 0.00001
[…]
MLP layer [2, 3] [2, 2] [2, 3] [2, 3]
[…]
$\lambda_{fm}$ 0.5 0.05 0.5 0.05
wd 0.01 0.00001 0.000001 0.00001
WuKong layer 8 4 8 4
Compr. dim 32 24 32 24
FMB size 200 200 200 200
[…]
Appendix B Supplementary Experiments
B.1 Efficiency Study
Empirical efficiency analysis. Empirically, we analysis the total training time of representative conventional CTR models and LLM-enhanced CTR models on Amazon Video Games and MovieLens-1M datasets with millions of interactions. As shown in Figure B.1, existing LLM-enhanced CTR methods (e.g., KAR, LLM-CF, CTRL and EASE) require over 290 times (on average) more computation time than conventional CTR models. With the increase in the number of interactions, this phenomenon will further exacerbate. In real-world industrial scenarios with interaction scales numbering in the billions, these LLM-enhanced CTR models will face more significant efficiency challenges.
(a) Video Games
(b) MovieLens-1M
Figure B.1: Empirical efficiency study on representative conventional CTR models and LLM-enhanced CTR models. The multiple increase in time cost is reported based on WuKong (the SOTA traditional CTR model).
Efficiency bottleneck analysis. The training time of all LLM-enhanced CTR methods can be roughly divided into four parts: fine-tuning, knowledge generation, semantic encoding and CTR prediction. To study the efficiency bottleneck of the four LLM-enhanced CTR baselines (KAR, LLM-CF, CTRL and EASE) and our LLaCTR, we conducted a detailed training time analysis on the Amazon Video Games and MovieLens-1M datasets.
Table B.3: Running time (s) details of LLM-enhanced CTR baselines and our LLaCTR.
[…]
LLM-CF 15,179 78,104 1,676 178 95,137
CTRL 0 0 5,839 329 6,168
EASE 32,221 0 1,747 125 34,093
[…]
LLM-CF 5,234 26,094 355 64 31,747
CTRL 0 0 1,401 141 1,542
EASE 12,813 0 49 43 12,905
[…]
The results are shown in Table B.3. The mark "/" represents that the corresponding phase does not exist in this method. We can derive:
  • 1) The efficiency bottleneck of KAR and LLM-CF is the knowledge generation. KAR requires to generate knowledge for a large number of users/items, while LLM-CF needs to generate the Chain-of-Thought (CoT) data on numerous instances (even if sampled as 10%). The reasoning and generation of LLMs consume a significant amount of time, and rise constantly as the length of generated text increases.
  • 2) The encoding time overhead of CTRL is significantly higher than other methods. Although encoding usually costs less time than Fine-tuning and Inference, CTRL still faces an efficiency bottleneck due to the large number of training instances that need to be encoded.
  • 3) The time overhead of fine-tuning with EASE is higher than other methods. EASE freezes the LLM and trains trains the semantic adaptor network during fine-tuning, but the frozen LLM still needs to be involved in the model forward phase. Besides, the parameter size of EASE’s semantic adaptor network (a BERT-Base [68] network with 12 transformer blocks and an additional 2 cross-attention layers) is about 110M, while the trainable parameter size of LoRA [89] based Llama3-8B [3] (used in LLM-CF and LLaCTR) is only approximately 3.4M (with the rank LoRA as 8), leading to a higher fine-tuning time overhead for EASE.
  • 4) The efficiency bottleneck of LLaCTR lies in fine-tuning. Although fine-tuning on the sampled field-feature data (in the SSFT module of LLaCTR) still incurs considerable time overhead compared to traditional CTR methods, it is much more efficient than fine-tuning on the numerous instance level or user/item level data (e.g., LLM-CF and EASE).
Time complexity analysis. To further study the training efficiency of LLaCTR and other LLM-enhanced CTR methods, we discuss the training time complexity of these models here. Let $A$ represent the average input token sequence length of the LLM, $B$ represent the average output token sequence length of the LLM, and $D$ represent the embedding dimension of the LLM. The average time overhead of the LLM during fine-tuning and encoding can be expressed as $H(A)$, while the average time overhead during generation can be expressed as $G(A,B)$. For the original CTR model, let $t$ represent the average time overhead of the model and $d$ represent the feature embedding dimension.
We use $U$ to represent the number of users, $I$ to represent the number of items, $N$ to represent the total number of instances, respectively. For LLaCTR, let $K$ represent the number of fields and $S$ represent the feature sample number of SSFT module. The time complexity of the SSFT module is $O(SKH(A)+KH(A))$, while the FRE module is $O(KDd)$ and the FRE module is $O(t+K^{2})$, respectively. The overall time complexity of LLaCTR is $O((S+1)KH(A)+KDd+(t+K^{2}))$. Since the time overhead on the LLM is much greater than the enhancement on CTR models, the SSFT module will dominate the overall time overhead, and the time complexity can be simplified to $O((S+1)KH(A))$. Similarly, the time complexity of KAR can be expressed as $O((U+I)G(A,B)+(U+I)H(A))$, the time complexity of LLM-CF can be expressed as $O(2NH(A)+NG(A,B))$, the time complexity of CTRL can be expressed as $O(NH(A))$, and the time complexity of EASE can be expressed as $O(2IH(A))$. Given that $K<SK\ll(U+I)<N$, LLaCTR exhibits significantly lower training time complexity compared to other LLM-enhanced CTR methods. This further highlights the efficiency of LLaCTR’s field-level enhancement approach.
B.2 Empirical Study
B.2.1 Fine-tuning Loss Study
To select the fine-tuning loss in the SSFT module of our LLaCTR, we compare the performance of language generative loss ($\mathcal{L}_{LG}$) with contrastive loss ($\mathcal{L}_{CL}$) on Amazon Gift Cards and Video Games datasets over DeepFM and WuKong backbones, and the results are shown in Table B.4.
Table B.4: Fine-tuning loss (in SSFT module) comparison between language generative loss ($\mathcal{L}_{LG}$) and contrastive loss ($\mathcal{L}_{CL}$).
[…]
LLaCTR ($\mathcal{L}_{LG}$) 0.4659 0.6794 4.73% 0.5755 0.6718 1.55%
[…]
LLaCTR ($\mathcal{L}_{LG}$) 0.4748 0.6800 4.14% 0.5795 0.6795 1.57%
[…]
We can find that the contrastive loss performs slightly better than the language generative loss, so we use the contrastive loss as the final fine-tuning loss in our experiments.
B.2.2 Alignment Loss Study
To select the feature representation enhancement loss in the FRE module of our LLaCTR, we compare the performance of mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$) on Amazon Gift Cards and Video Games datasets over DeepFM and WuKong backbones, and the results are shown in Table B.5.
Table B.5: Feature representation enhancement loss (in FRE module) comparison between mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$).
                    Method Gift Cards         Video Games
[…]
LLaCTR ($\mathcal{L}_{MSE}$) 0.4750 0.6767 3.15% 0.5760 0.6711 1.15%
LLaCTR ($\mathcal{L}_{CL}$) 0.4750 0.6763 2.92% 0.5762 0.6699 0.44%
[…]
LLaCTR ($\mathcal{L}_{MSE}$) 0.4755 0.6737 0.49% 0.5787 0.6795 1.57%
LLaCTR ($\mathcal{L}_{CL}$) 0.4753 0.6738 0.55% 0.5805 0.6788 1.17%
[…]
We can find that the KL divergence loss performs better than the MSE loss and CL loss, so we use the KL divergence loss as the final feature representation enhancement loss in our experiments.
B.3 Ablation Study
We conducted an ablation study on different datasets to study the contributions of each component of LLaCTR. For the Self-supervised Field-feature Fine-tuning (SSFT) module, we removed the entire instructing Fine-tuning component (w/o FT). For Feature Representation Enhancement (FRE) module and Feature Interaction Enhancement (FIE) module, we removed the entire alignment loss (w/o FRE) and the entire field interaction enhancement network (w/o FIE), respectively. The results are presented in Table B.6.
Table B.6: Ablation Study on LLaCTR.
[…]
w/o FIE 0.4667 0.6710 0.5767 0.6652 0.3391 0.7572 0.4019 0.8364
[…]
As can be seen, both the three components are important — removing SSFT, FRE or the FIE module would result in performance drops in most case. Delving deeper into the SSFT module, we observe that self-supervised field-feature fine-tuning on LLMs is indeed helpful. The field description encoded by the fine-tuned LLM has higher quality than that extracted from the pre-trained corpus. For FRE module, we observe that the developed KL-Divergence loss is also important. It can effectively inject semantic knowledge of fields into the feature representations. For FIE module, we find that for both explicit feature modeling and implicit feature modeling CTR methods, injecting field interaction scores can effectively guide feature interaction and bring AUC improvement.
B.4 Hyperparameters Sensitivity Study
Figure B.2 illustrates performance of LLaCTR with different hyper-parameters ($\lambda_{kl},\lambda_{fm}$) on Gift Cards and Video Games datasets, where $\lambda_{kl}$ and $\lambda_{fm}$ control the effects of Feature Representation Enhancement (FRE) and Feature Interaction Enhancement (FIE) respectively.
(a) $\lambda_{kl}$ on Gift Cards
(b) $\lambda_{kl}$ on Gift Cards
(c) $\lambda_{kl}$ on Gift Cards
(d) $\lambda_{fm}$ on Gift Cards
(e) $\lambda_{fm}$ on Gift Cards
(f) $\lambda_{fm}$ on Gift Cards
(g) $\lambda_{kl}$ on Video Games
(h) $\lambda_{kl}$ on Video Games
(i) $\lambda_{kl}$ on Video Games
(j) $\lambda_{fm}$ on Video Games
(k) $\lambda_{fm}$ on Video Games
(l) $\lambda_{fm}$ on Video Games
Figure B.2: Sensitivity analysis w.r.t. $\lambda_{kl}$, $\lambda_{fm}$.
We can observed the general trend is that the model’s performance would increase at the beginning and then drop as these parameters increase. This result validates the effectiveness of the Feature Representation Enhancement (FRE) module and Feature Interaction Enhancement (FIE) module. But over-emphasizing the introduced component would incur performance drops as it would relatively decline the contribution from original models. Finely tuning these hyper-parameters for best balance could achieve optimal performance.
B.5 Case Study
To visually inspect the semantic information captured by LLaCTR at the field-level interactions, we analyzed the field heatmap of LLaCTR on the Video Games and MovieLens-1M dataset, as shown in Figure B.3.
Figure B.3: Case study on the heatmap of LLM’s field embeddings.
We can observe several interesting results: 1) For the Video Games dataset, there is a significant heat value between the field embeddings of "average_rating" and "rating_number", which aligns with the common sense that "users tend to purchase items with both high average ratings and high sales numbers". Increasing the interaction weight between these fields can help the CTR model make better predictions. Additionally, "average_rating" and "rating_number" exhibit high heat values with several other fields such as "item_id" and "user_id", indicating the relative importance of these fields. This insight could potentially inspire research in feature selection. 2) For the MovieLens-1M dataset, the heat value between "UserID" and "MovieID" is relatively high, which to some extent explains the significant individual preference differences in MovieLens-1M. Injecting this information to guide the learning of feature interactions can assist the model in better capturing the semantic information in the dataset.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract accurately conclude this paper’s contribution, including the findings and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of this work is discussed in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper fully discloses all the information needed to reproduce the experimental results in Section 4 and Appendix A.5.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The data and code are provided in https://anonymous.4open.science/r/LLaCTR-EC46.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All the experimental details are included in Section 4 and Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The $p$-value of the main performance is reported in Table 1.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Sufficient information on the computer resources is provided in Figure 4 and Appendix B.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform in every respect with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This work proposed an effective and efficient LLM-enhanced CTR method and has no societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no risks for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The original paper that produced the codes and datasets are cited in the paper without any omissions.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The new assets introduced in the paper are well documented and the details of the code are included.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The LLM is used only for writing and editing in this paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 Field Matters: A Lightweight LLM-enhanced Method for CTR Prediction
-Abstract
+Abstract.
 Click-through rate (CTR) prediction is a fundamental task in modern recommender systems.
@@
 In our experiments, we integrate LLaCTR with six representative CTR models across four datasets, demonstrating its superior performance in terms of both effectiveness and efficiency compared to existing LLM-enhanced methods.
-Our code is available at https://anonymous.4open.science/r/LLaCTR-EC46.
-1 Introduction
-Click-through rate (CTR) prediction [1] is a core task in modern recommender systems (RS), with the objective of estimating the likelihood that a user will click on a specific item by modeling complex feature interactions.
-With the advent of Large Language Models (LLMs) and their remarkable abilities in content comprehension and semantic reasoning [2, 3], there has been a surge of interest in leveraging LLMs for CTR prediction.
-Recent studies mainly explored two paradigms: 1) LLMs as CTR predictors [4, 5, 6], where LLMs are either prompted or fine-tuned to directly perform CTR prediction; 2) LLM-enhanced CTR models [7, 8, 9, 10], where LLMs are leveraged to augment traditional CTR models by injecting semantic knowledge.
+Our code is available at https://github.com/istarryn/LLaCTR.
+Large language Model, CTR Prediction, Recommender Systems
+journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792387isbn: 979-8-4007-2307-0/2026/04submissionid: rfp1942ccs: Information systems Recommender systems
+1. Introduction
+Click-through rate (CTR) prediction (Yang and Zhai, 2022) is a core task in modern recommender systems (RS) (Chen et al., 2023; Zhang et al., 2025), with the objective of estimating the likelihood that a user will click on a specific item by modeling complex feature interactions.
+With the advent of Large Language Models (LLMs) and their remarkable abilities in content comprehension and semantic reasoning (Achiam et al., 2023; Dubey et al., 2024), there has been a surge of interest in leveraging LLMs for CTR prediction.
+Recent studies mainly explored two paradigms: 1) LLMs as CTR predictors (Geng et al., 2022; Bao et al., 2023b; Geng et al., 2024), where LLMs are either prompted or fine-tuned to directly perform CTR prediction; 2) LLM-enhanced CTR models (Li et al., 2023; Xi et al., 2024; Sun et al., 2024; Qiu et al., 2024), where LLMs are leveraged to augment traditional CTR models by injecting semantic knowledge.
 Despite their encouraging performance, these approaches encounter significant practical limitations, particularly regarding computational efficiency and economic feasibility.
@@
 Although the second paradigm alleviates inference latency by retaining conventional CTR models for online serving, it still suffers prohibitive training costs.
-Specifically, these methods typically operate at the instance level [7, 9, 11] or the user/item level [8, 12, 10], requiring LLMs to process detailed textual descriptions for large-scale data instances or user/item entities.
-Empirically, we find that existing LLM-enhanced CTR methods (e.g., KAR[8], LLM-CF[9], CTRL[7] and EASE [10]) require over 290 times (on average) more computation time than baseline models on typical Amazon Video Games and MovieLens-1M datasets with millions of interactions (Figure 4 and Appendix B.1).
-This computational burden becomes even more pronounced when scaling to industrial settings involving billions of instances.
+Specifically, these methods typically operate at the instance level (Li et al., 2023; Sun et al., 2024; Lin et al., 2024a) or the user/item level (Xi et al., 2024; Wang et al., 2024; Qiu et al., 2024), requiring LLMs to process detailed textual descriptions for large-scale data instances or user/item entities.
+Empirically, we find that existing LLM-enhanced CTR methods (e.g., KAR(Xi et al., 2024), LLM-CF(Sun et al., 2024), CTRL(Li et al., 2023) and EASE (Qiu et al., 2024)) require over 290 times (on average) more computation time than baseline models on typical Amazon Video Games and MovieLens-1M datasets with millions of interactions (cf.
+Figure 1).
 These challenges naturally motivate a critical research question: How can we leverage LLMs to enhance CTR models in a more efficient and cost-effective manner?
+Figure 1. Empirical efficiency study on representative conventional CTR models and LLM-enhanced CTR models. The multiple increase in time cost is reported based on WuKong.
 To address this challenge, we explore a novel field-level enhancement paradigm that utilizes LLMs to extract semantic knowledge from feature fields to enhance CTR models.
 As illustrated in Figure 2, fields denote categories of features in CTR prediction.
-This paradigm offers two principal advantages: 1) Computational Efficiency: The number of fields is orders of magnitude lower than that of instances or entities (e.g., thousands vs.
-billions in industrial RS), yielding a drastic reduction in LLM computational overhead.
+This paradigm offers two principal advantages: 1) Computational Efficiency: The number of fields is orders of magnitude lower than that of instances or entities (e.g., hundreds vs.
+millions in the real RS), yielding a drastic reduction in LLM computational overhead.
 2) Crucial Semantic Information: Conventional CTR methods primarily utilize field IDs as categorical indicators, often neglecting the rich semantic information inherent in field descriptions.
 In fact, field semantics can significantly enhance both feature representation and feature interaction modeling — the two core components contributing the success of CTR prediction.
-For example, recognizing that the feature “4.7” belongs to the field “average rating” provides insight into item quality; while understanding the relationship between fields such as “user income” and “item price” can inform the importance of feature interactions.
+For example, recognizing that the feature “4.7” belongs to the field “average rating” provides insight into item quality; understanding the relationship between fields such as “user income” and “item price” can inform the importance of feature interactions.
 Although field-level semantic knowledge may not be as rich as cumbersome instance-level semantic information, its conciseness and importance offer substantial potential for improving model performance.
-Motivated by these analyses, we introduce LLaCTR, a novel Lightweight LLM-enhanced CTR method through field-level enhancement. LLaCTR comprises two key components:
+Motivated by these analyses, we introduce LLaCTR, a novel Lightweight LLM-enhanced CTR method through field-level enhancement, comprising two key components:
   • Self-supervised Field-feature Fine-tuning.
@@
 Moreover, LLaCTR outperforms recent LLM-enhanced methods while incurring substantially lower computational overhead (by a factor of 10–100 times).
-Figure 1: Current LLM-enhanced CTR paradigm versus our Field-level enhancement paradigm.
-2 Preliminaries
-Task Formulation.
-This work focuses on click-through rate (CTR) prediction [13, 14], a core task in many personalized services (e.g., recommender systems).
+Overall, this work makes the following contributions:
+  • We identify the limitations of existing LLM-based CTR methods and advocate for leveraging field-level semantic knowledge to efficiently enhance traditional CTR models.
+  • We propose a novel LLM-enhanced CTR method, LLaCTR, which employs self-supervised fine-tuning to distill high-quality field knowledge from LLMs and integrates this knowledge to improve both feature representation and feature interaction modeling.
+  • We conduct extensive experiments to demonstrate that LLaCTR significantly improves the accuracy of traditional CTR models while being highly efficient and cost-effective.
+Figure 2. Current LLM-enhanced CTR paradigm versus our field-level enhancement paradigm.
+2. Preliminaries
+2.1. Task Formulation
+This work focuses on click-through rate (CTR) prediction (Lin et al., 2023; Wang et al., 2022), a core task in many personalized services (e.g., recommender systems).
 Let $\mathcal{D}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N}$ denote the training dataset of user-item historical interaction records, where $\mathbf{X}_{i}=[\mathbf{x}_{i1},\mathbf{x}_{i2},\ldots,\mathbf{x}_{iK}]$ represents the input features across $K$-field for the $i$-th instance, and $\mathbf{x}_{ik}$ denotes the feature of the $k$-th field for the instance.
@@
 The goal of CTR is to learn a model from $\mathcal{D}$ that can accurately predict the click probability of a new instance.
-Traditional CTR Models. Traditional CTR models [15, 16, 17, 18, 19, 20] primarily adopt deep learning paradigms and can generally be abstracted into three core components:
+2.2. Traditional CTR Models
+Traditional CTR models (Rendle, 2010; Guo et al., 2017; Pan et al., 2018; Sun et al., 2021; Mao et al., 2023b; Zhang et al., 2024) primarily adopt deep learning paradigms and can generally be abstracted into three core components:
 1) Feature Embedding Layer. Given the high-dimensional and sparse nature of raw input features, feature embedding is commonly employed to map the raw features into dense representations:
-$\mathbf{E}_{i}=\text{EmbeddingLayer}(\mathbf{X}_{i}),$ (1)
-where $\mathbf{E}_{i}=[\mathbf{e}_{i1},\mathbf{e}_{i2},\ldots,\mathbf{e}_{iK}]\in\mathbb{R}^{K\times D}$ represents the learned feature embeddings of the instance $i$, with $\mathbf{e}_{ik}$ denoting the D-dimensional embedding vector for $k$-th field feature.
+(1) $\mathbf{E}_{i}=\text{EmbeddingLayer}(\mathbf{X}_{i}),$
+where $\mathbf{E}_{i}=[\mathbf{e}_{i1},\mathbf{e}_{i2},\ldots,\mathbf{e}_{iK}]\in\mathbb{R}^{K\times D}$ represents the learned feature embeddings of the instance $\mathbf{X}_{i}$, with $\mathbf{e}_{ik}$ denoting the D-dimensional embedding vector for $k$-th field feature.
 2) Feature Interaction Layer. This crucial component is designed to effectively capture the complex feature interactions:
-$\Phi(\mathbf{X}_{i})=\text{FeatureInteraction}(\mathbf{X}_{i},\mathbf{E}_{i}),$ (2)
+(2) $\Phi(\mathbf{X}_{i})=\text{FeatureInteraction}(\mathbf{X}_{i},\mathbf{E}_{i}),$
 where $\Phi(\cdot)$ is the hidden representation learned from feature interactions.
-Among various architectures, the most representative one is the factorization machine (FM) [15], which explicitly model bi-level feature interactions:
-$\Phi_{FM}(\mathbf{X}_{i})={w_{0}}+\sum\limits_{k=1}^{K}{{\mathbf{w}_{k}}^{T}{\mathbf{x}_{ik}}}+\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}\text{>}}}})},$ (3)
-where $w_{0},\mathbf{w}_{k}$ denotes the learnable weights, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the inner product of two vectors; $\varsigma(.)$ denotes the sum of all elements in the matrix.
-For convenience, here we adopt vector notation to represent the formula of Factorization Machines (FM), which is mathematically equivalent to the original formulation presented in [15].
+Among various architectures, the most representative one is the factorization machine (FM) (Rendle, 2010), which explicitly model bi-level feature interactions:
+(3) $\small\Phi_{FM}(\mathbf{X}_{i})={w_{0}}+\sum\limits_{k=1}^{K}{{\mathbf{w}_{k}}^{T}{\mathbf{x}_{ik}}}+\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}\text{>}}}})},$
+where $w_{0},\mathbf{w}_{k}$ denotes the learnable weights, and $\cdot$$\cdot$> denotes the inner product of two vectors; $\varsigma(.)$ denotes the sum of all elements in the matrix.
+For convenience, here we adopt vector notation to represent the formula of Factorization Machines (FM), which is mathematically equivalent to the original formulation presented in (Rendle, 2010).
 The features are explicitly bi-level interacted with $\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}$ and its contribution is controlled by their embeddings $\text{<}\mathbf{e}_{ik},\mathbf{e}_{il}\text{>}$.
-Building on FM, recent works have explored various extensions [21, 22, 23], such as introducing field-aware matrices [24, 17, 18] or incorporating neural layers [16, 25].
-Meanwhile, other studies have proposed novel architectures, such as convolution operator [26, 27] and self-attention [28, 29], to better capture complex feature interactions.
+Building on FM, recent works have explored various extensions (Xiao et al., 2017; Lu et al., 2021; Tao et al., 2020), such as introducing field-aware matrices (Juan et al., 2016; Pan et al., 2018; Sun et al., 2021) or incorporating neural layers (Guo et al., 2017; Zhang et al., 2019).
+Meanwhile, other studies have proposed novel architectures, such as convolution operator (Liu et al., 2019a; Li et al., 2019) and self-attention (Zhou et al., 2018, 2019), to better capture complex feature interactions.
 3) Prediction Layer. Finally, the prediction layer transforms $\Phi(\mathbf{X}_{i})$ into the model prediction $\hat{y_{i}}$:
-$\hat{y_{i}}=\sigma[\text{PredictionLayer}(\Phi(\mathbf{X}_{i}))],$ (4)
-where $\sigma(\cdot)$ is the sigmoid function. The model is typically optimized with the binary cross-entropy (BCE) loss [30] with:
-$\mathcal{L}_{BCE}(y,\hat{y})=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\left(\hat{y_{i}}\right)+(1-y_{i})\log\left(1-\hat{y_{i}}\right)\right],$ (5)
-While these traditional CTR models have made significant progress in the past few decades, they are often plagued by the issue of semantic information loss.
-Specifically, Some methods have utilized field IDs as categorical indicators [24, 17, 18], but the rich semantic information inherent in field descriptions has not been exploited.
-This limitation motivates us to leverage LLMs to enrich the traditional CTR models with field-level knowledge.
-LLM-enhanced CTR Models.
-To fully exploit semantic knowledge, Large Language Models (LLMs) [2, 3] have been extensively investigated for enhancing traditional CTR models [8, 7, 9, 10, 31].
+(4) $\hat{y_{i}}=\sigma[\text{PredictionLayer}(\Phi(\mathbf{X}_{i}))],$
+where $\sigma(\cdot)$ is the sigmoid function. The model is typically optimized with the binary cross-entropy (BCE) loss (Mao et al., 2023a) with:
+[…]
+While these traditional CTR models have made significant progress in the past few decades, they often suffer from the semantic information loss.
+Specifically, some methods rely solely on field IDs as categorical indicators (Juan et al., 2016; Pan et al., 2018; Sun et al., 2021), neglecting the rich semantic information in field descriptions.
+This limitation motivates our use of LLMs to enhance traditional CTR models with field-level semantic knowledge.
+Figure 3. The overall framework of proposed LLaCTR.
+2.3. LLM-enhanced CTR Models
+To fully exploit semantic knowledge, LLMs have been extensively investigated for enhancing traditional CTR models.
 As illustrated in Figure 2, existing LLM-enhanced CTR methods primarily operate at the instance level or user/item level.
 Typically, these approaches first organize the features of users, items, or instances into textual descriptions, which are then provided as prompts to the LLM for reasoning or summarization.
-The resulting semantic knowledge, often at the user/item or instance level, is further encoded into semantic embeddings to augment traditional CTR models (see Section 5 for details).
+The resulting semantic knowledge, often at the user/item or instance level, is further encoded into semantic embeddings to augment traditional CTR models.
 However, due to the large scale of users/items and instances in real-world applications, such strategies incur prohibitive computational costs during training, inference and encoding.
-For example, on representative datasets such as Amazon Video Games and MovieLens-1M, the total training time of existing LLM-enhanced CTR methods (e.g., KAR [8], LLM-CF [9], CTRL [7] and EASE [10]) is over 290 times (on average) that of standard CTR models on average (see Appendix B.1 for details).
+For example, on representative datasets such as Amazon Video Games and MovieLens-1M, the total training time of existing LLM-enhanced CTR methods (e.g., KAR (Xi et al., 2024), LLM-CF (Sun et al., 2024), CTRL (Li et al., 2023) and EASE (Qiu et al., 2024)) is over 290 times (on average) that of standard CTR models on average (as shown in Figure 1).
 Moreover, the complex and voluminous semantic knowledge generated at the user/item or instance level is often difficult for traditional CTR models to effectively assimilate and utilize.
 Given these limitations, it is imperative to explore new paradigms for integrating semantic knowledge into CTR models.
-3 Methodology
-Figure 2: The overall framework of proposed LLaCTR.
-It contains 1) SSFT: leveraging self-supervised field-feature fine-tuning to improve LLMs’ ability to capture field semantics; 2) FRE: leveraging field semantic embeddings to guide the learning of the feature representation; 3) FIE: leveraging field semantic embeddings to enhance the models of feature interactions.
+3. Methodology
 In this section, we introduce LLaCTR, which enhances CTR prediction through field-level semantic knowledge from LLMs.
-As shown in Figure 2, LLaCTR first employs self-supervised field-feature fine-tuning to improve LLMs’ ability to capture field semantics (Section 3.1), and then leverage the field knowledge distilled from LLMs to enhance both feature representation (Section 3.2) and feature interactions (Section 3.3).
-3.1 Self-supervised Field-feature Fine-tuning (SSFT)
-LLMs are pre-trained on open-domain natural language corpora [2, 3], which often lack domain-specific knowledge relevant to CTR prediction.
+As shown in Figure 3, LLaCTR first employs self-supervised field-feature fine-tuning to improve LLMs’ ability to capture field semantics (SSFT), and then leverages the field knowledge distilled from LLMs to enhance both feature representation (FRE) and feature interactions (FIE).
+3.1. Self-supervised Field-feature Fine-tuning (SSFT)
+LLMs are pre-trained on open-domain natural language corpora (Achiam et al., 2023; Dubey et al., 2024), which often lack domain-specific knowledge relevant to CTR prediction.
 This limitation can hinder their ability to distill high-quality field semantic knowledge.
 Specifically, LLMs may not fully understand the meaning of field descriptions in CTR tasks or their relationships with features.
-To address this, we propose a self-supervised fine-tuning strategy that prompts LLMs to predict the field to which a given feature belongs, thereby enhancing their understanding of field semantic.
+To address this, we propose a self-supervised fine-tuning strategy that prompts LLMs to predict the field to which a given feature belongs, thereby enhancing their understanding of field semantics.
 Specifically, let $F_{k}$ represent the descriptions of the $k$-th field, and $f^{k}_{j}$ denote the description of the $j$-th feature within field $k$.
@@
 These prompt-response pairs are used to fine-tune LLMs with language generative loss:
-$\mathcal{L}_{LG}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log P_{LLM}\left(F_{k}\mid\mathcal{P}(f^{k}_{j})\right),$ (6)
+(6) $\mathcal{L}_{LG}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log P_{LLM}\left(F_{k}\mid\mathcal{P}(f^{k}_{j})\right),$
 where $P_{LLM}\left(F_{k}\mid\mathcal{P}(f^{k}_{j})\right)$ represents the LLM’s probability of generating the correct field description $F_{k}$ given the prompt $\mathcal{P}(f^{k}_{j})$.
 This loss improves the LLM’s ability to capture field-feature correlations, injecting domain knowledge and enabling the generation of higher-quality semantic representations.
-Besides the language generative loss, we find the contrastive loss [32] is also effective:
-$\mathcal{L}_{CL}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log\frac{\exp(cos(En(\mathcal{P}(f_{j}^{k})),En(F_{k}))/\tau)}{\sum_{l=1}^{K}\exp(cos(En(\mathcal{P}(f_{j}^{k})),En(F_{l}))/\tau)},$ (7)
-where $cos(\cdot)$ denotes the cosine similarity function, $\tau$ is the temperature parameter and $En(\cdot)$ are the LLM’s encodings of the language descriptions, respectively.
+Besides the language generative loss, we find the contrastive loss (He et al., 2020) is also effective:
+(7) $\small\mathcal{L}_{CL}(\mathcal{P}(f^{k}_{j}),F_{k};\theta)=-\log\frac{\exp(cos(\mathcal{E}(\mathcal{P}(f_{j}^{k})),\mathcal{E}(F_{k}))/\tau)}{\sum_{l}\exp(cos(\mathcal{E}(\mathcal{P}(f_{j}^{k})),\mathcal{E}(F_{l}))/\tau)},$
+where $cos(\cdot)$ denotes the cosine similarity function, $\tau$ is the temperature parameter and $\mathcal{E}(\cdot)$ are the LLM’s encodings of the language descriptions, respectively.
 The contrastive loss is designed to align the semantic embeddings of prompts with their corresponding answers, thereby effectively injecting feature-field correlation knowledge in embeddings.
-Since the contrastive loss operates directly on embeddings, which are subsequently used to enhance CTR models, this work prioritizes the use of contrastive loss and empirically find slightly better performance than generative loss (Appendix B.2).
-After fine-tuning, the LLM’s encodings of field descriptions are distilled:
-$\textbf{h}_{k}=En(F_{k}).$ (8)
+Since the contrastive loss operates directly on embeddings, which are subsequently used to enhance CTR models, this work prioritizes the use of contrastive loss and empirically find slightly better performance than generative loss (see Appendix B.2.1 for details).
+After fine-tuning, the LLM’s encodings of field descriptions are extracted:
+(8) $\textbf{h}_{k}=\mathcal{E}(F_{k}).$
 For convenience, we collect the field embeddings as a matrix $\textbf{H}=\{\textbf{h}_{k}\}_{k=1}^{K}$. These field embeddings will further utilized to enhance traditional CTR models.
-Figure 3: The template of self-supervised tuning.
+Figure 4. The template of self-supervised fine-tuning.
 Notably, such field-level operations are highly efficient, as the LLMs require to process only a small number of fields (typically less than 100) rather than millions of instances or user/item pairs in the experimental datasets.
-This difference in scale is even more pronounced in real-world industrial recommender systems, where there may be thousands of fields, but hundreds of millions of users/items and billions of instances.
 Some readers may be concerned about the potential inefficiency of self-supervised fine-tuning.
@@
 Empirical comparisons of efficiency can be found in Figure 5.
-The detailed time complexity analyses of LLaCTR and other LLM-enhanced baselines are provided in Appendix B.1.
-3.2 Feature Representation Enhancement (FRE)
+The detailed time efficiency analyses of LLaCTR and other LLM-enhanced baselines are provided in Section 4.2.2.
+3.2. Feature Representation Enhancement (FRE)
 The quality of feature representations is critical for the success of CTR models.
-However, existing methods often overlook the semantic knowledge contained in field descriptions.
+However, existing methods often overlook the semantic knowledge contained in field descriptions (Lin et al., 2024a; Xi et al., 2024).
 To address this, we transfer the field semantic knowledge distilled from LLMs to enhance feature embeddings.
 Given the embedding space gap between LLMs and CTR models, we first map original field embeddings $h_{k}$ into the embedding space of CTR models using a learnable adaptor:
-$\textbf{h}^{{}^{\prime}}_{k}=\text{Adaptor}(\textbf{h}_{k}),$ (9)
-where $\text{Adaptor}(\cdot)$ can be implemented as a simple linear transformation. We then align feature embeddings with their corresponding field embeddings through normalized KL-divergence [33]:
-$\mathbf{h}^{{}^{\prime}}_{k}\leftarrow\text{Softmax}(\mathbf{h}^{{}^{\prime}}_{k}),\ \mathbf{e}_{ik}\leftarrow\text{Softmax}(\mathbf{e}_{ik}),$ (10)
-$\mathcal{L}_{KL}=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{D}_{\text{KL}}(\mathbf{e}_{ik},\mathbf{h}^{{}^{\prime}}_{k})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{e}_{ik}\cdot\ln({\mathbf{e}_{ik}}/{\mathbf{h}_{k}^{{}^{\prime}}}).$ (11)
+(9) $\textbf{h}^{{}^{\prime}}_{k}=\text{Adaptor}(\textbf{h}_{k}),$
+where $\text{Adaptor}(\cdot)$ can be implemented as a simple linear transformation.
+We then align feature embeddings with their corresponding field embeddings through normalized KL-divergence (Tian et al., 2019):
+(10) $\mathbf{h}^{{}^{\prime}}_{k}\leftarrow\text{Softmax}(\mathbf{h}^{{}^{\prime}}_{k}),\ \mathbf{e}_{ik}\leftarrow\text{Softmax}(\mathbf{e}_{ik}),$
+(11) $\small\mathcal{L}_{KL}=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{D}_{\text{KL}}(\mathbf{e}_{ik},\mathbf{h}^{{}^{\prime}}_{k})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{e}_{ik}\cdot\ln({\mathbf{e}_{ik}}/{\mathbf{h}_{k}^{{}^{\prime}}}).$
 Specifically, here we employ normalized KL-divergence, as it eliminates the influence of embedding magnitude and focuses on aligning the underlying distributions.
-We also experimented with other commonly used alignment losses, such as contrastive loss, but found that normalized KL-divergence consistently yields superior performance (Appendix B.2).
+We also experimented with other commonly used alignment losses, such as contrastive loss, but found that normalized KL-divergence consistently yields superior performance (see Appendix B.2.2 for details).
 By minimizing $\mathcal{L}_{KL}$, feature embeddings are able to absorb semantic knowledge from the field embeddings, resulting in higher-quality representations.
@@
 The overall training objective combines semantic alignment with the original CTR objective:
-$\mathcal{L}=\mathcal{L}_{BCE}+\lambda_{kl}\mathcal{L}_{KL},$ (12)
-where $\lambda_{kl}$ serves as a hyperparameter balancing the contributions of the two components.
-3.3 Feature Interaction Enhancement (FIE)
-Field information has been shown beneficial to improve feature interaction modeling in CTR tasks [24, 15, 18].
-However, these traditional method only utilize field ID information by employing implicit embedding layers [24] or learnable parameters matrices [15, 18], without taking the textual information of fields into account.
+(12) $\mathcal{L}=\mathcal{L}_{BCE}+\lambda_{kl}\mathcal{L}_{KL},$
+where hyperparameter $\lambda_{kl}$ balances their contributions.
+3.3. Feature Interaction Enhancement (FIE)
+Field information has been shown beneficial to improve feature interaction modeling in CTR tasks (Juan et al., 2016; Rendle, 2010; Sun et al., 2021).
+However, these traditional method only utilize field ID information by employing implicit embedding layers (Juan et al., 2016) or learnable parameters matrices (Rendle, 2010; Sun et al., 2021), without taking the textual information of fields into account.
 Field semantic information provides new insights into feature interactions.
@@
 To leverage this insight, we first compute the importance of feature interactions of two fields based on their semantic embeddings:
-$m_{ij}^{F}=\text{FieldInteraction}(\mathbf{h}_{i},\mathbf{h}_{j}),$ (13)
-[…]
+(13) $m_{ij}^{F}=\text{FieldInteraction}(\mathbf{h}_{i},\mathbf{h}_{j}),$
+where $\text{FieldInteraction}(\cdot)$ denotes the field interaction layer, which can be implemented by various architecture.
+In practice, we find that cosine similarity is effective — intuitively, the fields with similar semantic may provide more insights on prediction.
+For convenience, we collect all pairwise field interaction scores into a matrix $\mathbf{M}^{F}$.
+Subsequently, we employ a learnable adaptor to re-scale these importance scores, i.e., $\mathbf{M}^{{}^{\prime}}=\text{Adaptor}(\mathbf{M}^{F})$, where the adaptor can be simply implemented via a linear layer.
 The field-aware importance scores derived from semantic embeddings are then utilized to guide the learning of feature interactions.
 Specifically, we explicitly incorporate these scores into a bi-level feature interaction modeling, thereby enhancing the feature interaction layer of CTR methods:
-$\Phi^{new}(\mathbf{X}_{i})=\Phi(\mathbf{X}_{i})+\lambda_{fm}\cdot\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}^{T}\text{>}m_{kl}^{{}^{\prime}}}}})},$ (14)
+(14) $\small\Phi^{new}(\mathbf{X}_{i})=\Phi(\mathbf{X}_{i})+\lambda_{fm}\cdot\sum\limits_{k=1}^{K}{\sum\limits_{l=k+1}^{K}\varsigma({{{\mathbf{x}_{ik}{\mathbf{x}_{il}}^{T}\text{<}\textbf{e}_{ik},{\textbf{e}_{il}}^{T}\text{>}m_{kl}^{{}^{\prime}}}}})},$
 where the field-aware importance score $m_{kl}^{{}^{\prime}}$ modulates the strength of feature interactions.
@@
 In our experiments, we incorporate our LLaCTR module into six representative CTR methods, and observe substantial performance improvements across almost all cases.
-Detailed results are presented in Section 4, and the implementation details are provided in the Appendix A.4.
-4 Experiments
-4.1 Experimental Setup
-Datasets.
-We adopt four widely-used public datasets in LLM-enhanced CTR task: Amazon Gift Cards, Amazon Video Games, Amazon Digital Music and MovieLens-1M.
-We closely refer to [11, 8] and partition each dataset into training, validation, and testing sets with ratios of 8:1:1 according to the timestamp.
-The details of the above datasets are summarized in Appendix A.1.
-Metrics.
-Following previous work [8, 34, 35], we employed two widely-used metrics LogLoss (binary cross-entropy loss) and AUC (Area under the ROC curve) to evaluate performance.
-RelaImpr [36], the relative AUC improvement compared to the basic CTR model [28, 7] is also reported.
-The readers may refer to Appendix A.2 for more details.
-Baselines.
-For comparisons, we selected KAR (RecSys’24) [8], LLM-CF (CIKM’24) [9], CTRL (TORS’23) [7] and EASE (CIKM’24) [10] as LLM-enhanced baselines.
-These baselines are closely related and representative LLM-enhanced CTR methods.
-For fair comparisons, unless otherwise specified, we consistently used the Llama3-8B [3] as the LLM.
-We also include two representative PLM-enhanced methods - CELA (arXiv’24) [12] and ClickPrompt (WWW’24) [11] for comparisons.
-We integrated these baseline methods into the following representative traditional CTR models: 1) Classic FM-based methods: FM (ICDM’10) [15] and DeepFM (IJCAI’17) [16]; 2) field-based CTR methods: FwFM (WWW’18) [17] and FmFM (WWW’21) [18]; 3) The state-of-the-art CTR methods: FinalMLP (AAAI’23) [19] and WuKong (ICML’24) [20].
-The readers may refer to Appendix A.4 for more details about these methods.
-Hyperparameter settings.
-A grid search is utilized to find the optimal hyperparameters.
-For all compared methods, we closely refer to the configurations provided in their respective publications to ensure their optimal performance.
-For LLaCTR, $\lambda_{kl}$ and $\lambda_{fm}$ are tuned in {0.01, 0.05, 0.1, 0.3, 0.5, 0.7, 1}, and the feature sampling number in fine-tuning is set at 1000.
-The detailed hyperparameter settings are provided in Appendix A.5.
-4.2 Experimental Results
-In this section, we primarily focus on performance comparison, efficiency analysis, and ablation studies.
-Additional experimental results, including hyperparameter sensitivity analyses and case studies, are provided in Appendix B.
-Table 1: Performance comparisons of LLaCTR with existing PLM/LLM-enhanced methods on six CTR backbone models.
+Detailed results are presented in Section 4, and the implementation details are provided in the Appendix A.
+Table 1. Statistics of the datasets.
+  Dataset Gift Cards Video Games Digital Music MovieLens-1M
+#Field 13 13 13 8
+#Feature 1,981,330 60,119,995 1,695,642 9,001,881
+#User 132,732 2,766,656 100,952 6,040
+#Item 1,137 137,249 70,511 3,706
+#Interaction 152,410 4,624,615 130,434 1,000,209
+Density $1.01\times 10^{-3}$ $1.22\times 10^{-5}$ $1.83\times 10^{-5}$ $4.47\times 10^{-2}$
+Table 2.
+Performance comparisons of LLaCTR with existing PLM/LLM-enhanced methods.
 The results of PLM-enhanced methods are reported on their best backbones.
 The best result is bolded and the blue-colored zone indicates that LLaCTR is better than the basic CTR backbone.
-The line "Average RelaImpr" indicates the average RelaImpr of LLaCTR over six CTR backbones.
-The mark "*" indicates the improvement is statistically significant ($p$-value $<0.05$).
-[…]
-[…]
-[…]
-[…]
-[…]
+The mark "*" indicates the improvement is statistically significant ($p$-value <0.05<0.05).
+  Gift Cards Video Games Digital Music       MovieLens-1M
+Method Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr
+CELA 0.4658 0.6746 1.02% 0.5798 0.6775 0.44% 0.3364 0.7638 0.37% 0.4031 0.8373 0.24%
+ClickPrompt 0.4656 0.6754 1.48% 0.5747 0.6799 1.79% 0.3356 0.7652 0.90% 0.4025 0.8379 0.42%
+Base 0.4606 0.6621 0.5805 0.6673 0.3359 0.7612 0.4081 0.8354
 KAR 0.4551 0.6643 1.34% 0.5803 0.6705 1.93% 0.3410 0.7607 -0.21% 0.4073 0.8358 0.13%
@@
 EASE 0.4534 0.6667 2.84% 0.5804 0.6701 1.69% 0.3359 0.7625 0.48% 0.4101 0.8357 0.10%
-[…]
-[…]
+FM LLaCTR 0.4500 0.6673 3.21%* 0.5801 0.6718 2.70%* 0.3362 0.7621 0.32% 0.4050 0.8364 0.30%*
+Base 0.4751 0.6713 0.5762 0.6692 0.3478 0.7582 0.4190 0.8342
 KAR 0.4753 0.6778 3.80% 0.5871 0.6720 1.65% 0.3439 0.7587 0.18% 0.4207 0.8335 -0.20%
@@
 EASE 0.4748 0.6739 1.52% 0.5807 0.6738 2.74% 0.3360 0.7604 0.83% 0.4180 0.8349 0.22%
-[…]
-[…]
+DeepFM LLaCTR 0.4657 0.6797 4.90%* 0.5834 0.6744 3.10%* 0.3350 0.7687 4.06%* 0.4049 0.8354 0.38%*
+Base 0.4664 0.6691 0.5762 0.6683 0.3420 0.7553 0.4017 0.8365
 KAR 0.4661 0.6694 0.19% 0.5791 0.6741 3.42% 0.3432 0.7575 0.88% 0.4020 0.8360 -0.15%
@@
 EASE 0.4661 0.6703 0.74% 0.5795 0.6765 4.84% 0.3415 0.7571 0.71% 0.4030 0.8369 0.12%
-[…]
-[…]
+FwFM LLaCTR 0.4659 0.6718 1.61%* 0.5827 0.6780 5.73%* 0.3362 0.7585 1.27%* 0.4014 0.8367 0.05%
+Base 0.4579 0.6747 0.5824 0.6706 0.3315 0.7628 0.4077 0.8325
 KAR 0.4595 0.6754 0.42% 0.5767 0.6711 0.29% 0.3347 0.7640 0.43% 0.4062 0.8332 0.21%
@@
 EASE 0.4590 0.6753 0.33% 0.5750 0.6729 1.35% 0.3311 0.7637 0.33% 0.4056 0.8334 0.27%
-[…]
-[…]
+FmFM LLaCTR 0.4560 0.6761 0.78%* 0.5725 0.6765 3.45%* 0.3323 0.7615 -0.50% 0.4055 0.8335 0.30%*
+Base 0.4774 0.6720 0.5830 0.6761 0.3423 0.7557 0.4086 0.8301
 KAR 0.4881 0.6726 0.35% 0.5820 0.6817 3.19% 0.3429 0.7546 -0.41% 0.4094 0.8290 -0.33%
@@
 EASE 0.4773 0.6736 0.93% 0.5776 0.6796 1.97% 0.3414 0.7560 0.12% 0.4049 0.8309 0.24%
-[…]
-[…]
+FinalMLP LLaCTR 0.4717 0.6818 5.71%* 0.5738 0.6832 4.02%* 0.3411 0.7552 -0.19% 0.4009 0.8340 1.18%*
+Base 0.4755 0.6728 0.5819 0.6767 0.3402 0.7514 0.4088 0.8323
 KAR 0.4794 0.6721 -0.43% 0.5927 0.6758 -0.53% 0.3429 0.7506 -0.31% 0.4099 0.8316 -0.21%
@@
 EASE 0.4748 0.6758 1.71% 0.5783 0.6801 1.91% 0.3381 0.7556 1.67% 0.4031 0.8334 0.33%
-[…]
-[…]
+WuKong LLaCTR 0.4745 0.6802 4.24%* 0.5763 0.6819 2.91%* 0.3377 0.7564 1.99%* 0.4027 0.8397 2.21%*
+Average RelaImpr 3.41% 3.65% 1.16% 0.74%
+4. Experiments
+We aim to answer the following research questions:
+  • $\mathbf{RQ1:}$ How does LLaCTR perform compared with the state-of-the-art CTR methods?
+  • $\mathbf{RQ2:}$ How is the efficiency of LLaCTR compared with existing LLM-enhanced CTR methods?
+  • $\mathbf{RQ3:}$ What are the impacts of different components of LLaCTR?
+4.1. Experimental Setup
+Figure 5. AUC and training time of compared methods. The “sam-num” is the sampling feature number of each field, and "w/o FT" represents fine-tuning has been removed.
+4.1.1. Datasets
+Following previous work (Lin et al., 2024a; Xi et al., 2024), we conduct experiments on four widely used public datasets in LLM-enhanced CTR task: Amazon Gift Cards, Amazon Video Games, Amazon Digital Music and MovieLens-1M.
+  • Amazon Review Datasets [^1] https://amazon-reviews-2023.github.io/ are well-known e-commercial datasets (Sun et al., 2024; Li et al., 2023; Lin et al., 2024a; Bao et al., 2023a) with ratings ranging from 1 to 5.
+Following (Lin et al., 2024a), we binarize the ratings with a threshold of 4 and use the 5-core setting, i.e., all users and items have at least 5 interactions.
+  • MovieLens-1M Dataset [^2] https://grouplens.org/datasets/MovieLens/1m/ is a benchmark movie recommendation dataset from Movielens with ratings ranging from 1 to 5.
+We binarize the ratings with a threshold of 4, while removing neutral samples with ratings equal to 3 following (Li et al., 2023).
+The dataset statistics are summarized in Table 1.
+Following (Li et al., 2023; Lin et al., 2024a), we organize the review behaviors in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1 after preprocessing.
+4.1.2. Evaluation Metrics
+Following previous work (Lin et al., 2024a; Song et al., 2019), we employ two widely-used metrics LogLoss (binary cross-entropy loss) and AUC (area under the ROC curve) to evaluate performance.
+The RelaImpr (Yan et al., 2014) (relative AUC improvement) is also reported.
+Notably, slightly higher AUC or lower LogLoss (e.g., 0.1%) can be regarded as significant improvement in CTR prediction, as indicated by previous studies (Huang et al., 2019; Song et al., 2019).
+4.1.3. Baselines
+For comparisons, we selected KAR (RecSys’24) (Xi et al., 2024), LLM-CF (CIKM’24) (Sun et al., 2024), CTRL (TORS’23) (Li et al., 2023) and EASE (CIKM’24) (Qiu et al., 2024) as LLM-enhanced baselines.
+These baselines are closely related and representative LLM-enhanced CTR methods.
+We also include two representative PLM-enhanced methods — CELA (arXiv’24) (Wang et al., 2024) and ClickPrompt (WWW’24) (Lin et al., 2024a) for comparisons.
+We integrated these baseline methods into the following representative traditional CTR models: 1) Classic FM-based methods: FM (ICDM’10) (Rendle, 2010) and DeepFM (IJCAI’17) (Guo et al., 2017); 2) Field-based CTR methods: FwFM (WWW’18) (Pan et al., 2018) and FmFM (WWW’21) (Sun et al., 2021); 3) The state-of-the-art CTR methods: FinalMLP (AAAI’23) (Mao et al., 2023b) and WuKong (ICML’24) (Zhang et al., 2024).
+The readers may refer to Appendix A for more details about these methods.
+4.1.4. Implementation Details
+We use the Adam (Kingma and Ba, 2014) optimizer to train all the CTR models.
+The learning rate (lr) is set as 0.001 and the weight decay (wd) is tuned in {1e-2, 1e-3, 1e-4, 1e-5, 1e-6, 0}.
+We fix the feature embedding size to 32 and batch size to 4096 for all the backbones and datasets.
+For all compared methods, we closely refer to the configurations provided in their respective publications to ensure their optimal performance.
+For LLaCTR fine-tuning, we fix the batch size as 128, the learning rate of LLM as 1e-4 and rank of LoRA (Hu et al., 2021) as 8, respectively.
+During LLaCTR enhancement, the $\lambda_{kl}$ and $\lambda_{fm}$ are tuned in {0.01, 0.05, 0.1, 0.3, 0.5, 0.7, 1}.
+We set the feature sampling number as 1000 by default, and set temperature parameter $\tau$ of contrastive loss into 0.02 following previous work (Wang et al., 2023).
+All methods are implemented in PyTorch and run on 8 Nvidia A800 GPUs.
+More implementation details are provided in the Appendix A.
+4.2. Experimental Results
+In this section, we focus on performance comparison, efficiency analysis, and ablation studies.
+Additional results, including time complexity analysis, hyperparameter sensitivity analysis and more details, are provided in Appendix B.
+4.2.1. Performance Comparison (RQ1)
+Table 2 shows the performance comparison of the proposed LLaCTR against the baseline methods. We observe that:
 Comparing with traditional CTR models.
-From the Table 1, we observe LLaCTR can be applied to various types of basic CTR models, and yield performance gains in most cases.
+LLaCTR can be applied to various types of basic CTR models, and yield performance gains in most cases.
 The improvements brought by LLaCTR are impressive, achieving an average AUC relative improvement of 2.24% over all basic CTR models across four datasets.
@@
 Comparing with PLM/LLM-enhanced CTR models.
-We observe that LLaCTR outperformed two representative PLM-enhanced CTR models (CELA and ClickPrompt) and four LLM-enhanced CTR models (KAR, LLM-CF, CTRL and EASE) in most cases (89%), which validates the effectiveness of our field-level enhancement paradigm.
-This result is highly surprising, as we just use the lightweight field knowledge rather than the cumbersome instance-level (or user/item-level) knowledge utilized by recent work.
-The reason may be : 1) field knowledge is indeed important and helpful for CTR prediction; 2) the lightweight field knowledge has been effectively exploited, enhancing both feature representation and feature interactions.
-Efficiency Analysis.
+We observe that LLaCTR outperforms two representative PLM-enhanced CTR models (CELA and ClickPrompt) and four LLM-enhanced CTR models (KAR, LLM-CF, CTRL and EASE) in most cases (89%), which validates the effectiveness of our field-level enhancement paradigm.
+This result is highly surprising, as we only utilize the lightweight field knowledge rather than the cumbersome instance-level (or user/item-level) knowledge utilized by recent work.
+The reasons may be 1) field knowledge is indeed important and helpful for CTR prediction; 2) the lightweight field knowledge has been effectively exploited, enhancing both feature representation and feature interactions.
+4.2.2. Efficiency Analysis (RQ2)
+In this section, we conduct an in-depth analysis on the training efficiency of different LLM-enhanced CTR methods.
+In terms of inference efficiency, since all these methods pre-store LLM knowledge and ultimately depend on CTR models for recommendations, the variation is much smaller than that observed during training (cf.
+Table 8 in Appendix B.3).
 The AUC performance and the total training time of compared methods are shown in Figure 5.
-The performance of LLM-enhanced methods are reported on their best backbones.
+The performances of LLM-enhanced methods are reported on their best backbones.
 We observe: 1) Compared to three representative LLM-enhanced methods, LLaCTR can improve efficiency by 10-100x and achieve better performance.
-2) With the increase of sampling number in the SSFT module, both the training time and AUC of LLaCTR will increase.
-This indicates increasing the number of samples used for self-supervised learning can yield higher-quality field semantic knowledge, but require more training time cost.
+2) With the increase of the sampling number in the SSFT module, both the training time and AUC of LLaCTR will increase.
+This indicates increasing the number of samples used for self-supervised learning can yield higher-quality field semantic knowledge, but require more training time.
 3) Even without fine-tuning, LLaCTR still achieves decent results.
 This suggests our LLaCTR could be applied in resource-constrained scenarios, enhancing existing CTR methods with incurring limited additional time cost.
-Ablation Study.
-We further conduct the ablation study, where the Self-supervised Fine-tuning (SSFT) module, Feature Representation Enhancement (FRE) module or Feature Interaction Enhancement (FIE) module has been removed respectively.
+Besides, we also study the efficiency bottleneck and the time complexity of these LLM-enhanced methods, the readers may refer to Appendix B.1 for more details.
+4.2.3. Ablation Study (RQ3)
+We further conduct the ablation study, where the Self-supervised Fine-tuning (SSFT) module, Feature Representation Enhancement (FRE) module or Feature Interaction Enhancement (FIE) module is removed respectively.
 The results are presented in Table 3.
-As can be seen, both the three components are important — with few exceptions, removing SSFT (w/o FT), FRE (w/o FRE) or FIE ($\lambda_{fm}$) module would result in performance drops.
-Figure 4: AUC and training time of compared methods. The “sam-num” represents the sampling feature number for each field of LLaCTR, and "w/o FT" represents fine-tuning has been removed.
-Table 2: Ablation study on LLaCTR with DeepFM and WuKong backbones. The results on other backbones are presented in Appendix B.3.
-[…]
-[…]
-[…]
-w/o FT 0.4687 0.6759 0.5771 0.6732 0.3393 0.7683 0.4064 0.8352
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-5 Related Work
-Traditional CTR Prediction.
-CTR Prediction serves as a core function module in personalized online services [37, 1, 38, 39, 16, 40].
-The key idea of traditional CTR prediction models is to capture the feature interaction patterns, which indicates the combination relationships of multiple features.
-The feature interaction layer is generally implemented by a combination of various operations.
-Product operators include classical bi-level models [15, 41, 42] and higher-order [43, 44, 45, 46, 47] models.
-Convolutional operators [48, 27, 26] utilize CNN [49] and GCN [50] to capture the implicit high order interaction.
-Due to the success of attention mechanism, attention-based methods [51, 21, 34, 35, 28, 29] also achieved great success.
-In recent years, several field-aware CTR methods have been proposed [24, 17, 18], however, these methods relied solely on field IDs as categorical indicators without leveraging the rich semantic information inherent in field descriptions.
-LLMs as CTR predictors.
-Initial efforts explored the usage of pre-trained language models as CTR predictors by reformulating recommendation tasks as natural language [4, 52, 53].
-With the advent of LLMs and their remarkable abilities in content comprehension and semantic reasoning [2, 3], recent efforts have shifted to explore using larger-scale Language Models to generate CTR prediction results [54, 55, 56, 57, 58, 59, 60, 31], e.g., TALLRec [5] uses instruction tuning with user’s historical interactions, enabling LLMs to achieve decent CTR prediction capability in zero-shot and few-shot scenarios, while HLLM [61] utilizes two independent LLMs to model users and items separately, and jointly trains them with both generative and discriminative recommendation losses.
-However, the high inference latency and resource demands of LLMs poses significant obstacles to their deployment in real-world large-scale recommender systems.
-Although recent studies have proposed various acceleration strategies, e.g., BAHE [6] freezes the shallow layers of LLMs to encode users’ atomic interactions and pre-stores them offline to reduce the number of parameters involved in training and inference, while Rella [62] reduces inference overhead by shortening user interaction sequences through the elimination of items that are semantically irrelevant to the target item, the latency and resource consumption nonetheless remain substantial due to the LLMs’ large parameter counts and complex architectures.
-LLM-enhanced CTR Prediction.
-To address the efficiency challenges inherent in LLMs-based CTR predictors, the LLM-enhanced CTR paradigm has been extensively investigated as an alternative approach.
-This paradigm primarily focus on leveraging the powerful semantic encoding and knowledge reasoning capabilities of LLMs to enhance traditional CTR models [63, 8, 9, 7, 10].
-For example, KAR [8] leverages the LLM as a open-word knowledge base to enhance the profile of users and items.
-LLM-CF [9] uses LLMs to generate a chains of thought (CoT) for sampled instances, which is further retrieved to enhance CTR models.
-CTRL [7] utilizes the LLM as the feature encoder for instances and uses a fine-grained contrastive learning scheme to align the knowledge.
-EASE [10] trains a semantic adaptor to align item embeddings with LLM, and use it for encoding during inference.
-Other work also focuses on specific CTR scenarios, such as cross-domain[64, 65] and interpretability [66, 67].
+As can be seen, both the three components are important — removing SSFT, FRE or the FIE module would result in performance drops in most case.
+Delving deeper into the SSFT module, we observe that self-supervised field-feature fine-tuning on LLMs is indeed helpful.
+The field description encoded by the fine-tuned LLM has higher quality than that extracted from the pre-trained corpus.
+For FRE module, we observe that the developed KL-Divergence loss is also important.
+It can effectively inject semantic knowledge of fields into the feature representations.
+For FIE module, we find that for both explicit feature modeling and implicit feature modeling CTR methods, injecting field interaction scores can effectively guide feature interaction and bring AUC improvement.
+Table 3. Ablation study on DeepFM and WuKong. Results on the other backbones are shown in Appendix B.4.
+                         Method Gift Games Music Movie
+DeepFM Base 0.6713 0.6692 0.7582 0.8342
+w/o FT 0.6759 0.6732 0.7683 0.8352
+w/o $\lambda_{kl}$ 0.6764 0.6699 0.7582 0.8343
+w/o $\lambda_{fm}$ 0.6787 0.6741 0.7685 0.8351
+LLaCTR 0.6797 0.6744 0.7687 0.8354
+WuKong Base 0.6728 0.6767 0.7514 0.8323
+w/o FT 0.6800 0.6777 0.7558 0.8322
+w/o $\lambda_{kl}$ 0.6730 0.6784 0.7524 0.8332
+w/o $\lambda_{fm}$ 0.6763 0.6783 0.7494 0.8347
+LLaCTR 0.6802 0.6819 0.7564 0.8397
+5. Related Work
+5.1. Traditional CTR Prediction
+CTR prediction is a core functional module in personalized online services, where the key idea is to capture feature interaction patterns that capture the combinational relationships among multiple features.
+Traditional methods employ various operations for feature interaction, including product-based operators (Cheng et al., 2016; Wang et al., 2021), convolutional networks (Liu et al., 2015), and attention mechanisms (Huang et al., 2019).
+Recently, field-aware CTR methods (Pan et al., 2018; Sun et al., 2021) have gained attention, though they primarily rely on field IDs as categorical indicators without fully leveraging the semantic richness of field descriptions.
+5.2. LLMs as CTR predictors
+Initial efforts explored the pre-trained language models as CTR predictors by reformulating recommendation tasks as NLP problems (Geng et al., 2022; Cui et al., 2022; Liu et al., 2022).
+With the advent of LLMs in content comprehension and semantic reasoning (Achiam et al., 2023; Dubey et al., 2024), LLMs have demonstrated strong potential in CTR prediction (Bao et al., 2023b; Chen et al., 2024).
+However, LLMs suffer from high inference latency and resource demands.
+While acceleration techniques exist (e.g., BAHE (Geng et al., 2024) for parameter reduction, Rella (Lin et al., 2024b) for sequence shortening), their computational overhead remains prohibitive for real-world deployment due to massive model sizes.
+5.3. LLM-enhanced CTR Prediction
+To address the efficiency challenges inherent in LLM-based CTR predictors, the LLM-enhanced CTR paradigm has been extensively investigated as an alternative approach.
+This paradigm primarily focuses on leveraging the powerful semantic encoding and knowledge reasoning capabilities of LLMs to enhance traditional CTR models (Li et al., 2023; Qiu et al., 2024).
+For example, KAR (Xi et al., 2024) leverages LLMs to enhance the knowledge of users and items.
+LLM-CF (Sun et al., 2024) uses LLMs to generate chains of thought (CoT) for enhancement.
+CTRL (Li et al., 2023) utilizes the LLM as the feature encoder for instances and uses contrastive learning to align the knowledge.
+EASE (Qiu et al., 2024) trains a semantic adaptor to align item embeddings with LLM.
+Other work also focuses on specific CTR scenarios, such as cross-domain(Fu et al., 2023) and interpretability (Yu et al., 2024).
 Despite their promising performance, they involve enhancements at the instance-level or user/item-level, which results in great inefficiency issue in dealing with large-scale instances or entities in practice.
 Our approach is applied on the more efficient and cost-effective field-level, without incurring superabundant computational overhead.
-There are also two related topics: 1) PLM-enhanced CTR methods: Early studies have leveraged pretrained language models (PLMs, like BERT [68] and RoBERTa[69]) to enhance CTR models via encoding textual features [70, 71, 72, 73, 53, 11, 12], e.g., CELA [12] pretrains a PLM on item descriptions and aligns item textual embeddings with ID-based embeddings for enhancement.
-ClickPrompt [11] uses a CTR model to encode instances into soft tokens, which are then treated as prefix tokens in a PLM for generating instance-level semantic knowledge.
-However, the capacity of early PLMs is generally inferior to that of recent LLMs, and recent studies have demonstrated the superior performance of LLM-enhanced CTR methods.
-Consequently, our work aims to further investigate and advance the state-of-the-art in LLM-enhanced CTR modeling.
-2) LLM-enhanced recommendation: Recent years also witnessed the integration of LLMs on other recommendation tasks,e.g., sequential recommendation [74, 75, 59, 76, 77, 78], collaborative filtering [79, 80, 65, 81], cross-domain recommendation [82, 83].
-It is important to note that CTR prediction differs from these tasks in that it typically involves various features and emphasizes modeling feature interactions.
-Thus, our LLaCTR is specifically designed to leverage LLMs for improved feature representation and interactions in CTR prediction.
-6 Conclusion and Limitations
-This work proposes a novel lightweight method, LLaCTR, which leverages LLMs to extract semantic knowledge from feature fields for enhancing CTR models.
+5.4. Other Related Work
+There are also two related topics: 1) PLM-enhanced CTR methods: Early studies have leveraged pretrained language models (PLMs, such as BERT (Devlin et al., 2019)) to enhance CTR models, e.g., CELA (Wang et al., 2024) pretrains PLMs on item features and aligns them with ID embeddings for enhancement.
+ClickPrompt (Lin et al., 2024a) uses the ID embeddings as prefix soft tokens in PLMs for generating instance-level knowledge.
+However, the capacity of early PLMs is generally inferior to that of recent LLMs, and recent studies have demonstrated the superior performance of LLM-enhanced methods.
+Consequently, our work aims to further investigate and advance the LLM-enhanced CTR modeling.
+2) LLM-enhanced recommendation: Recent years have also witnessed the integration of LLMs on other recommendation tasks (Cui et al., 2025; Wang et al., 2025a, 2026), e.g., sequential recommendation (Cui et al., 2024; Wang et al., 2025b), collaborative filtering (Ren et al., 2024; Lin et al., 2025).
+Notably, CTR prediction differs from these tasks in that it typically involves various features and emphasizes modeling feature interactions.
+Thus, our LLaCTR is specifically designed to improve feature representation and interactions in CTR prediction.
+6. Conclusion
+This work proposes a lightweight method, LLaCTR, which leverages LLMs to extract semantic knowledge from feature fields for enhancing CTR models.
 Specifically, LLaCTR employs a self-supervised field-feature fine-tuning to distill high-quality field-level knowledge, which is subsequently utilized to improve both feature representation and feature interaction modeling.
 Extensive experiments demonstrate the superiority of LLaCTR in both predictive accuracy and computational efficiency.
-Nevertheless, a current limitation of LLaCTR is that it leverages field-level semantic knowledge to guide bi-level feature interaction modeling.
-Exploring more advanced strategies for incorporating field knowledge to capture higher-order feature interactions remains an interesting and promising direction for future research.
-Appendix A Experimental Details
-A.1 Datasets
-We conduct experiments on four real-world datasets from different recommendation scenarios. The datasets statistics are summarized in Table A.1.
-Table A.1: Statistics of the datasets.
-[…]
-[…]
-#Feature 1,981,330 60,119,995 1,695,642 9,001,881
-#User 132,732 2,766,656 100,952 6,040
-#Item 1,137 137,249 70,511 3,706
-#Interaction 152,410 4,624,615 130,434 1,000,209
-Density $1.01\times 10^{-3}$ $1.22\times 10^{-5}$ $1.83\times 10^{-5}$ $4.47\times 10^{-2}$bigstrut[b]
-  • Amazon Review Datasets [^1] https://amazon-reviews-2023.github.io/ are widely-used e-commercial datasets [9, 7, 11, 84] from Amazon website with ratings ranging from 1 to 5.
-We use the Gift Cards, Video Games and Digital Music subsets in our experiments.
-Following the standard practice in Lin et al.
-[11], we binarize the ratings with a threshold of 4 and use 5-core setting, i.e., all users and items have at least 5 interactions.
-  • MovieLens-1M Dataset [^2] https://grouplens.org/datasets/MovieLens/1m/ is a movie recommendation dataset from Movielens website with ratings ranging from 1 to 5.
-We binarize the ratings with a threshold of 4, while removing neutral samples with ratings equal to 3 following previous work [7].
-Following [11], we organized the review behaviors in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1 after preprocessing.
-All the experiments are implemented with PyTorch and run on 4x Nvidia A800 GPUs.
-A.2 Metrics
-To evaluate the performance of CTR prediction methods, we closely follow [28, 8, 11, 7] and adopt AUC (area under the ROC curve), LogLoss (binary cross-entropy loss) and RelaImpr (AUC relative improvement) as the evaluation metrics:
-  • AUC [85]: AUC is the area under the ROC curve and measures the goodness of order by ranking all the items with predicted CTR. For a random guesser, the value of AUC is 0.5.
-  • LogLoss: Logloss is the logarithmic loss value of binary cross-entropy loss and quantifies the accuracy of predicted probabilities by penalizing false classifications.
-  • RelaImpr [36]: RelaImpr measures the AUC relative improvement over basic CTR models and is defined as below:
-    $RelaImpr=(\frac{\text{AUC (measured\ model)}-0.5}{\text{AUC (base\ model)}-0.5}-1)\times 100\%.$ (15)
-Slightly higher AUC or lower Log Loss (e.g., 0.001) can be regarded as significant improvement in CTR prediction, as indicated by previous studies [34, 35].
-A.3 Baselines
+In the future, it would be valuable to investigate more sophisticated strategies for integrating field semantics to further enhance CTR methods.
+Acknowledgements.
+This work is supported by the Zhejiang Province “JianBingLingYan+X” Research and Development Plan (2025C02020).
+We thank the reviewers for their valuable and insightful suggestions that improve the paper.
+Appendix A Appendix A: Experimental Details
+A.1. Baselines
 We reproduced the following LLM-enhanced CTR methods as baselines in our experiments:
-  • KAR (RecSys’24) [8]: KAR leverages the LLM as a open-word knowledge base to enhance the profile of users and items, and then pass the encoded user/item embeddings through a knowledge adaptation network to enhance traditional CTR models.
-  • LLM-CF (CIKM’24) [9]: LLM-CF uses LLMs to generate a base of chains of thought (CoT) for sampled instances, which are further retrieved to enhance CTR models.
-Consistent with the original paper, we sampled 10% of the instances as sampled data for fine-tuning and CoT data generation, and during downstream CTR prediction, we also retrieve the Top 4 samples closest in semantics from the CoT dataset as enhanced knowledge.
-  • CTRL (TORS’23) [7]: CTRL utilizes the LLM as the feature encoder for instances and uses a fine-grained contrastive learning scheme to align cross-modal knowledge.
-In our implementation, we used Llama3-8B [3] as the LLM to encode instances embeddings for fair comparison.
-  • EASE (CIKM’24) [10]: EASE trains a semantic adaptor network to align item embeddings with the LLM, and then use it for encoding during inference.
-Following the original paper, we inited the semantic adapter using BERT-Base [68] with 12 transformer blocks, and cross-attention layers are added to the 6-th and 12-th transformer blocks.
-For fair comparisons, unless otherwise specified, we consistently used the Llama3-8B [3] as the LLM in all the LLM-enhanced CTR methods.
+  • KAR (RecSys’24) (Xi et al., 2024): KAR leverages the LLM as a open-word knowledge base to enhance the profile of users and items, and then pass the encoded user/item embeddings through a knowledge adaptation network to enhance traditional CTR models.
+  • LLM-CF (CIKM’24) (Sun et al., 2024): LLM-CF uses LLMs to generate a base of chains of thought (CoT) for sampled instances, which are further retrieved to enhance CTR models.
+Consistent with the original paper, we sampled 10% of the instances as sampled data for fine-tuning and CoT data generation.
+  • CTRL (TORS’23) (Li et al., 2023): CTRL utilizes the LLM as the feature encoder for instances and uses a fine-grained contrastive learning scheme to align cross-modal knowledge.
+  • EASE (CIKM’24) (Qiu et al., 2024): EASE trains a semantic adaptor network to align item embeddings with the LLM, and then use it for encoding during inference.
+Following the original paper, we inited the semantic adapter using BERT-Base (Devlin et al., 2019) with 12 transformer blocks, and cross-attention layers are added to the 6-th and 12-th transformer blocks.
+For fair comparisons, unless otherwise specified, we consistently used the Llama3-8B (Dubey et al., 2024) as the LLM in all the LLM-enhanced CTR methods.
 Besides, we also reproduced two representative PLM-enhanced methods:
-  • CELA (arXiv’24) [12]: CELA first pretrains a PLM on item descriptions, then aligns item textual embeddings with pretrained ID-based embeddings, and finally merges them into CTR models for enhancement.
-We selected Robert-base [69] as the PLM as it performs the best in the original paper.
-  • ClickPrompt (WWW’24) [11]: ClickPrompt uses a CTR model to encode instances into soft tokens, which are then treated as prefix tokens in a PLM for generating instance-level semantic knowledge.
-Following the original paper, we selected Robert-base [69] as the PLM in our experiment.
-A.4 Backbones
+  • CELA (arXiv’24) (Wang et al., 2024): CELA first pretrains a PLM on item descriptions, then aligns item textual embeddings with pretrained ID-based embeddings, and finally merges them into CTR models for enhancement.
+We selected Robert-base (Liu et al., 2019b) as the PLM as it performs the best in the original paper.
+  • ClickPrompt (WWW’24) (Lin et al., 2024a): ClickPrompt uses a CTR model to encode instances into soft tokens, which are then treated as prefix tokens in a PLM for generating instance-level semantic knowledge.
+Following the original paper, we selected Robert-base (Liu et al., 2019b) as the PLM in our experiment.
+A.2. Backbones
+For all the backbone models, we reused the implement of an open-source CTR prediction library [^3] https://github.com/reczoo/FuxiCTR - FuxiCTR (Zhu et al., 2022), following the previous work (Mao et al., 2023b; Zhu et al., 2023; Li et al., 2024).
 We selected six representative CTR methods as backbones in our experiments, including:
-  • FM (ICDM’10) [15]: FM introduces factorization machines to model pairwise feature interactions via factorized parameters.
-  • DeepFM (IJCAI’17) [16]: DeepFM combines factorization machines with deep neural networks to jointly learn low-order and high-order feature interactions.
+  • FM (ICDM’10) (Rendle, 2010): FM introduces factorization machines to model pairwise feature interactions via factorized parameters.
+  • DeepFM (IJCAI’17) (Guo et al., 2017): DeepFM combines factorization machines with deep neural networks to jointly learn low-order and high-order feature interactions.
 We fix the DNN layers at [300, 300, 128] in our experiments.
-  • FwFM (WWW’18) [17]: FwFM proposes field-weighted factorization machines that learn field-specific interaction weights to better capture heterogeneous feature relationships.
-  • FmFM (WWW’21) [18]: FmFM enhances FM with field-matrix factorization using learnable projection matrices per field pair, enabling more expressive cross-field interactions.
-  • FinalMLP (AAAI’23) [19]: FinalMLP designs two Stream Feature Interaction networks to progressively refine feature representations learned by MLP.
-We select the MLP hidden units size from {400, 500} and the MLP layer number from {2, 3}, respectively.
-  • WuKong (ICML’24) [20]: WuKong employs an interaction stack on the WuKong layer to capture feature interactions, where each WuKong layer consists of a Factorization Machine Block (FMB) and a Linear Compress Block.
-We select the interaction layer number from {4, 8}, the FMB size from {128, 200}, the FMB layer from {2, 3} and the compression dimension from {24, 32} respectively.
+  • FwFM (WWW’18) (Pan et al., 2018): FwFM proposes field-weighted factorization machines that learn field-specific interaction weights to better capture heterogeneous feature relationships.
+  • FmFM (WWW’21) (Sun et al., 2021): FmFM enhances FM with field-matrix factorization using learnable projection matrices per field pair, enabling more expressive cross-field interactions.
+  • FinalMLP (AAAI’23) (Mao et al., 2023b): FinalMLP designs two Stream Feature Interaction networks to progressively refine feature representations learned by MLP.
+We select the MLP hidden units size from {400, 500} and the MLP layer number from {2, 3}.
+  • WuKong (ICML’24) (Zhang et al., 2024): WuKong employs an interaction stack on the WuKong layer to capture feature interactions, where each WuKong layer consists of a Factorization Machine Block (FMB) and a Linear Compress Block.
+We select the interaction layer number from {4, 8}, the FMB layer from {2, 3} and the compression dimension from {24, 32} respectively.
 For the implementation of LLaCTR, the SSFT module and FRE module remained consistent across all backbones.
@@
   • 2) For implicit feature interaction backbones (FinalMLP and WuKong), the FIE module was treated as an additional plugin network to explicitly influence the second-order feature interactions learning and generate CTR prediction logit, which was later fused into the final prediction result of the original CTR model.
-The hyperparameter settings of each method are detailed in Appendix A.5.
-For all the backbone models, we reused the implement of an open-source CTR prediction library [^3] https://github.com/reczoo/FuxiCTR - FuxiCTR [86], following the previous work [19, 87, 47].
-A.5 Hyperparameters
-A.5.1 Hyperparameter Settings
-We use Adam [88] optimizer to train all the CTR models.
-The learning rate (lr) is set as 0.001 and the weight decay (wd) is tuned in {1e-2, 1e-3, 1e-4, 1e-5, 1e-6, 0}.
-We fix the feature embedding size to 32 and batch size to 4096 for all the backbones and datasets.
-For LLaCTR fine-tuning, we fix the batch size as 128, the learning rate of LLM as 1e-4 and rank of LoRA [89] as 8, respectively.
-We set the feature sampling number as 1000 by default, and set temperature parameter $\tau$ of contrastive loss into 0.02 following previous work [90].
-During LLaCTR enhancement, the $\lambda_{kl}$ and $\lambda_{fm}$ are tuned in {0.01, 0.05, 0.1, 0.3, 0.5, 0.7, 1}.
-A.5.2 Optimal Hyperparameters
-The hyperparameters we searched include: the $\lambda_{kl}$ for Feature Interaction Enhancement (FRE) module, the $\lambda_{fm}$ for Feature Interaction Enhancement (FIE) module, the weight decay (wd) of Adam [88] optimizer and other hyperparameters for backbone CTR models.
-Table A.2 shows the optimal hyperparameters, including four datasets (Gift Cards, Video Games, Digital Music and MovieLens-1M) and six backbones (FM, DeepFM, FwFM, FmFM, FinalMLP and WuKong).
-Table A.2: Optimal hyperparameters for 4 datasets and 6 backbones.
-[…]
-[…]
-[…]
-$\lambda_{fm}$ 0.01 0.01 0.7 0.5
-[…]
-[…]
-$\lambda_{fm}$ 0.5 0.5 0.3 0.7
-[…]
-[…]
-$\lambda_{fm}$ 0.1 0.1 0.3 0.3
-[…]
-[…]
-$\lambda_{fm}$ 0.1 0.1 0.05 0.7
-[…]
-[…]
-$\lambda_{fm}$ 0.1 0.3 0.01 0.05
-wd 0.01 0.00001 0.000001 0.00001
-[…]
-MLP layer [2, 3] [2, 2] [2, 3] [2, 3]
-[…]
-$\lambda_{fm}$ 0.5 0.05 0.5 0.05
-wd 0.01 0.00001 0.000001 0.00001
-WuKong layer 8 4 8 4
-Compr. dim 32 24 32 24
-FMB size 200 200 200 200
-[…]
-Appendix B Supplementary Experiments
-B.1 Efficiency Study
-Empirical efficiency analysis.
-Empirically, we analysis the total training time of representative conventional CTR models and LLM-enhanced CTR models on Amazon Video Games and MovieLens-1M datasets with millions of interactions.
-As shown in Figure B.1, existing LLM-enhanced CTR methods (e.g., KAR, LLM-CF, CTRL and EASE) require over 290 times (on average) more computation time than conventional CTR models.
-With the increase in the number of interactions, this phenomenon will further exacerbate.
-In real-world industrial scenarios with interaction scales numbering in the billions, these LLM-enhanced CTR models will face more significant efficiency challenges.
-(a) Video Games
-(b) MovieLens-1M
-Figure B.1: Empirical efficiency study on representative conventional CTR models and LLM-enhanced CTR models.
-The multiple increase in time cost is reported based on WuKong (the SOTA traditional CTR model).
-Efficiency bottleneck analysis.
-The training time of all LLM-enhanced CTR methods can be roughly divided into four parts: fine-tuning, knowledge generation, semantic encoding and CTR prediction.
-To study the efficiency bottleneck of the four LLM-enhanced CTR baselines (KAR, LLM-CF, CTRL and EASE) and our LLaCTR, we conducted a detailed training time analysis on the Amazon Video Games and MovieLens-1M datasets.
-Table B.3: Running time (s) details of LLM-enhanced CTR baselines and our LLaCTR.
-[…]
-[…]
-[…]
-LLM-CF 15,179 78,104 1,676 178 95,137
-CTRL 0 0 5,839 329 6,168
-EASE 32,221 0 1,747 125 34,093
-[…]
-[…]
-LLM-CF 5,234 26,094 355 64 31,747
-CTRL 0 0 1,401 141 1,542
-EASE 12,813 0 49 43 12,905
-[…]
-The results are shown in Table B.3. The mark "/" represents that the corresponding phase does not exist in this method. We can derive:
+Table 4. Fine-tuning loss (in SSFT module) comparison between language generative loss ($\mathcal{L}_{LG}$) and contrastive loss ($\mathcal{L}_{CL}$).
+                    Method Gift Cards         Video Games
+Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr
+DeepFM Base 0.4751 0.6713 0.5762 0.6692
+$\mathcal{L}_{LG}$ 0.4659 0.6794 4.73% 0.5755 0.6718 1.55%
+$\mathcal{L}_{CL}$ 0.4657 0.6797 4.90% 0.5834 0.6744 3.10%
+WuKong Base 0.4755 0.6728 0.5819 0.6767
+$\mathcal{L}_{LG}$ 0.4748 0.6800 4.14% 0.5795 0.6795 1.57%
+$\mathcal{L}_{CL}$ 0.4745 0.6802 4.24% 0.5763 0.6819 2.91%
+Table 5.
+Feature representation enhancement loss (in FRE module) comparison between mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$).
+                    Method Gift Cards         Video Games
+Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr
+DeepFM Base 0.4751 0.6713 0.5762 0.6692
+$\mathcal{L}_{MSE}$ 0.4750 0.6767 3.15% 0.5760 0.6711 1.15%
+$\mathcal{L}_{CL}$ 0.4750 0.6763 2.92% 0.5762 0.6699 0.44%
+$\mathcal{L}_{KL}$ 0.4657 0.6797 4.90% 0.5834 0.6744 3.10%
+WuKong Base 0.4755 0.6728 0.5819 0.6767
+$\mathcal{L}_{MSE}$ 0.4755 0.6737 0.49% 0.5787 0.6795 1.57%
+$\mathcal{L}_{CL}$ 0.4753 0.6738 0.55% 0.5805 0.6788 1.17%
+$\mathcal{L}_{KL}$ 0.4745 0.6802 4.24% 0.5763 0.6819 2.91%
+Appendix B Appendix B: Supplementary Experiments
+B.1. Efficiency Study
+B.1.1. Efficiency Bottleneck Analysis
+The training time of all LLM-enhanced CTR methods can be divided into four parts: fine-tuning, knowledge generation, semantic encoding and CTR prediction.
+To study the efficiency bottleneck, we conducted a detailed training time analysis on Video Games and MovieLens-1M datasets.
+The results are shown in Table 6.
+We can derive:
   • 1) The efficiency bottleneck of KAR and LLM-CF is the knowledge generation.
-KAR requires to generate knowledge for a large number of users/items, while LLM-CF needs to generate the Chain-of-Thought (CoT) data on numerous instances (even if sampled as 10%).
+KAR requires to generate knowledge for a large number of users/items, while LLM-CF needs to generate the CoT data on numerous instances (even if sampled as 10%).
 The reasoning and generation of LLMs consume a significant amount of time, and rise constantly as the length of generated text increases.
@@
 EASE freezes the LLM and trains trains the semantic adaptor network during fine-tuning, but the frozen LLM still needs to be involved in the model forward phase.
-Besides, the parameter size of EASE’s semantic adaptor network (a BERT-Base [68] network with 12 transformer blocks and an additional 2 cross-attention layers) is about 110M, while the trainable parameter size of LoRA [89] based Llama3-8B [3] (used in LLM-CF and LLaCTR) is only approximately 3.4M (with the rank LoRA as 8), leading to a higher fine-tuning time overhead for EASE.
+Besides, the parameter size of EASE’s semantic adaptor network (a BERT-Base (Devlin et al., 2019) network with 12 transformer blocks and an additional 2 cross-attention layers) is about 110M, while the trainable parameter size of LoRA (Hu et al., 2021) based Llama3-8B (Dubey et al., 2024) (used in LLM-CF and LLaCTR) is only approximately 3.4M (with the rank LoRA as 8), leading to a higher fine-tuning time overhead for EASE.
   • 4) The efficiency bottleneck of LLaCTR lies in fine-tuning.
 Although fine-tuning on the sampled field-feature data (in the SSFT module of LLaCTR) still incurs considerable time overhead compared to traditional CTR methods, it is much more efficient than fine-tuning on the numerous instance level or user/item level data (e.g., LLM-CF and EASE).
-Time complexity analysis.
+Table 6. Time Cost (s) details of LLM-enhanced CTR baselines and our LLaCTR.
+  Dataset Method        Time Cost in Details
+Fine-tuning Generation Enoding Prediction
+Video Games KAR 0 77,861 3,886 120
+LLM-CF 15,179 78,104 1,676 178
+CTRL 0 0 5,839 329
+EASE 32,221 0 1,747 125
+LLaCTR 599 0 12 124
+MovieLens-1M KAR 0 2,057 108 42
+LLM-CF 5,234 26,094 355 64
+CTRL 0 0 1,401 141
+EASE 12,813 0 49 43
+LLaCTR 415 0 11 34
+B.1.2. Time Complexity Analysis
 To further study the training efficiency of LLaCTR and other LLM-enhanced CTR methods, we discuss the training time complexity of these models here.
-Let $A$ represent the average input token sequence length of the LLM, $B$ represent the average output token sequence length of the LLM, and $D$ represent the embedding dimension of the LLM.
-The average time overhead of the LLM during fine-tuning and encoding can be expressed as $H(A)$, while the average time overhead during generation can be expressed as $G(A,B)$.
-For the original CTR model, let $t$ represent the average time overhead of the model and $d$ represent the feature embedding dimension.
-We use $U$ to represent the number of users, $I$ to represent the number of items, $N$ to represent the total number of instances, respectively.
-For LLaCTR, let $K$ represent the number of fields and $S$ represent the feature sample number of SSFT module.
+For the LLMs, let $A$ and $B$ denote the average input and output token sequence lengths of the LLM, respectively, and $D$ its embedding dimension.
+The average time overhead of the LLM can be expressed as $H(A)$ for fine-tuning/encoding and $G(A,B)$ for generation.
+For the original CTR model, $t$ denotes average time overhead and $d$ the feature embedding dimension.
+$U$, II, and $N$ denote the number of users, items, and total instances, respectively.
+In LLaCTR, let $K$ represent the number of fields and $S$ represent the feature sample number of SSFT module.
 The time complexity of the SSFT module is $O(SKH(A)+KH(A))$, while the FRE module is $O(KDd)$ and the FRE module is $O(t+K^{2})$, respectively.
@@
 Since the time overhead on the LLM is much greater than the enhancement on CTR models, the SSFT module will dominate the overall time overhead, and the time complexity can be simplified to $O((S+1)KH(A))$.
-Similarly, the time complexity of KAR can be expressed as $O((U+I)G(A,B)+(U+I)H(A))$, the time complexity of LLM-CF can be expressed as $O(2NH(A)+NG(A,B))$, the time complexity of CTRL can be expressed as $O(NH(A))$, and the time complexity of EASE can be expressed as $O(2IH(A))$.
+Similarly, the time complexity of KAR is $O((U+I)G(A,B)+(U+I)H(A))$, LLM-CF is $O(2NH(A)+NG(A,B))$, CTRL is $O(NH(A))$, and EASE is $O(2IH(A))$.
 Given that $K<SK\ll(U+I)<N$, LLaCTR exhibits significantly lower training time complexity compared to other LLM-enhanced CTR methods.
 This further highlights the efficiency of LLaCTR’s field-level enhancement approach.
-B.2 Empirical Study
-B.2.1 Fine-tuning Loss Study
-To select the fine-tuning loss in the SSFT module of our LLaCTR, we compare the performance of language generative loss ($\mathcal{L}_{LG}$) with contrastive loss ($\mathcal{L}_{CL}$) on Amazon Gift Cards and Video Games datasets over DeepFM and WuKong backbones, and the results are shown in Table B.4.
-Table B.4: Fine-tuning loss (in SSFT module) comparison between language generative loss ($\mathcal{L}_{LG}$) and contrastive loss ($\mathcal{L}_{CL}$).
-[…]
-[…]
-[…]
-LLaCTR ($\mathcal{L}_{LG}$) 0.4659 0.6794 4.73% 0.5755 0.6718 1.55%
-[…]
-[…]
-LLaCTR ($\mathcal{L}_{LG}$) 0.4748 0.6800 4.14% 0.5795 0.6795 1.57%
-[…]
-We can find that the contrastive loss performs slightly better than the language generative loss, so we use the contrastive loss as the final fine-tuning loss in our experiments.
-B.2.2 Alignment Loss Study
-To select the feature representation enhancement loss in the FRE module of our LLaCTR, we compare the performance of mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$) on Amazon Gift Cards and Video Games datasets over DeepFM and WuKong backbones, and the results are shown in Table B.5.
-Table B.5: Feature representation enhancement loss (in FRE module) comparison between mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$).
-                    Method Gift Cards         Video Games
-[…]
-[…]
-LLaCTR ($\mathcal{L}_{MSE}$) 0.4750 0.6767 3.15% 0.5760 0.6711 1.15%
-LLaCTR ($\mathcal{L}_{CL}$) 0.4750 0.6763 2.92% 0.5762 0.6699 0.44%
-[…]
-[…]
-LLaCTR ($\mathcal{L}_{MSE}$) 0.4755 0.6737 0.49% 0.5787 0.6795 1.57%
-LLaCTR ($\mathcal{L}_{CL}$) 0.4753 0.6738 0.55% 0.5805 0.6788 1.17%
-[…]
-We can find that the KL divergence loss performs better than the MSE loss and CL loss, so we use the KL divergence loss as the final feature representation enhancement loss in our experiments.
-B.3 Ablation Study
-We conducted an ablation study on different datasets to study the contributions of each component of LLaCTR.
-For the Self-supervised Field-feature Fine-tuning (SSFT) module, we removed the entire instructing Fine-tuning component (w/o FT).
-For Feature Representation Enhancement (FRE) module and Feature Interaction Enhancement (FIE) module, we removed the entire alignment loss (w/o FRE) and the entire field interaction enhancement network (w/o FIE), respectively.
-The results are presented in Table B.6.
-Table B.6: Ablation Study on LLaCTR.
-[…]
-[…]
-[…]
-[…]
-[…]
-w/o FIE 0.4667 0.6710 0.5767 0.6652 0.3391 0.7572 0.4019 0.8364
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-As can be seen, both the three components are important — removing SSFT, FRE or the FIE module would result in performance drops in most case.
-Delving deeper into the SSFT module, we observe that self-supervised field-feature fine-tuning on LLMs is indeed helpful.
-The field description encoded by the fine-tuned LLM has higher quality than that extracted from the pre-trained corpus.
-For FRE module, we observe that the developed KL-Divergence loss is also important.
-It can effectively inject semantic knowledge of fields into the feature representations.
-For FIE module, we find that for both explicit feature modeling and implicit feature modeling CTR methods, injecting field interaction scores can effectively guide feature interaction and bring AUC improvement.
-B.4 Hyperparameters Sensitivity Study
-Figure B.2 illustrates performance of LLaCTR with different hyper-parameters ($\lambda_{kl},\lambda_{fm}$) on Gift Cards and Video Games datasets, where $\lambda_{kl}$ and $\lambda_{fm}$ control the effects of Feature Representation Enhancement (FRE) and Feature Interaction Enhancement (FIE) respectively.
-(a) $\lambda_{kl}$ on Gift Cards
-(b) $\lambda_{kl}$ on Gift Cards
-(c) $\lambda_{kl}$ on Gift Cards
-(d) $\lambda_{fm}$ on Gift Cards
-(e) $\lambda_{fm}$ on Gift Cards
-(f) $\lambda_{fm}$ on Gift Cards
-(g) $\lambda_{kl}$ on Video Games
-(h) $\lambda_{kl}$ on Video Games
-(i) $\lambda_{kl}$ on Video Games
-(j) $\lambda_{fm}$ on Video Games
-(k) $\lambda_{fm}$ on Video Games
-(l) $\lambda_{fm}$ on Video Games
-Figure B.2: Sensitivity analysis w.r.t. $\lambda_{kl}$, $\lambda_{fm}$.
+B.2. Empirical Study
+B.2.1. Fine-tuning Loss Study
+For the fine-tuning loss in the SSFT module, we compare the performance of language generative loss ($\mathcal{L}_{LG}$) with contrastive loss ($\mathcal{L}_{CL}$) on over DeepFM and WuKong backbones, and the results are shown in Table 4.
+We observe that the contrastive loss performs slightly better than the language generative loss, so we use the contrastive loss as the final fine-tuning loss.
+Table 7. Ablation Study on the other backbones.
+                    Method Gift Games Music Movie
+FwFM Base 0.6691 0.6683 0.7553 0.8365
+w/o FT 0.6714 0.6757 0.7587 0.8365
+w/o FRE 0.6692 0.6729 0.7553 0.8366
+w/o FIE 0.6710 0.6652 0.7572 0.8364
+LLaCTR 0.6718 0.6780 0.7585 0.8367
+FmFM Base 0.6747 0.6706 0.7628 0.8325
+w/o FT 0.6759 0.6759 0.7614 0.8334
+w/o FRE 0.6750 0.6714 0.7608 0.8334
+w/o FIE 0.6753 0.6710 0.7612 0.8332
+LLaCTR 0.6761 0.6765 0.7615 0.8335
+FinalMLP Base 0.6723 0.6761 0.7557 0.8260
+w/o FT 0.6815 0.6787 0.7554 0.8330
+w/o FRE 0.6757 0.6772 0.7564 0.8332
+w/o FIE 0.6800 0.6775 0.7555 0.8324
+LLaCTR 0.6818 0.6832 0.7552 0.8340
+B.2.2. Alignment Loss Study
+For the feature representation enhancement loss in the FRE module, we compare the performance of mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$) over DeepFM and WuKong backbones, and the results are shown in Table 5.
+We find that the KL divergence loss performs better than the MSE loss and CL loss, which is more suitable for enhancement loss.
+B.3. Inference Efficiency Study
+Table 8. Inference Time (s) on LLM-enhanced CTR methods.
+  Dataset WuKong KAR LLM-CF CTRL EASE LLaCTR
+Video Games 1.57 1.66 1.84 1.81 1.73 1.63
+MovieLens-1M 0.32 0.41 0.53 0.46 0.45 0.41
+We compare the total inference time of different LLM-enhanced methods on the Video Games and MovieLens-1M datasets, with the results shown in Table 8 (WuKong serves as the backbone model).
+It can be observed that the differences among various LLM-enhanced methods during the inference stage are negligible and do not introduce significant latency compared to the backbone model.
+This is because these LLM-enhanced methods have pre-stored the knowledge of LLMs, requiring only lightweight networks and a small number of parameters to inject the knowledge into the inference process of traditional CTR models, thereby avoiding excessive additional latency.
+B.4. Ablation Study
+The ablation study results on the FwFM, FmFM and FinalMLP backbones are presented in Table 7.
+As can be seen, the results are consistent with those obtained from other backbones.
+Both the three modules are important, removing each would result in performance drops.
+B.5. Hyperparameters Sensitivity
+Figure 6 illustrates performance of LLaCTR with different hyper-parameters, where $\lambda_{kl}$ and $\lambda_{fm}$ control the effects of Feature Representation Enhancement (FRE) and Feature Interaction Enhancement (FIE) respectively.
 We can observed the general trend is that the model’s performance would increase at the beginning and then drop as these parameters increase.
@@
 Finely tuning these hyper-parameters for best balance could achieve optimal performance.
-B.5 Case Study
-To visually inspect the semantic information captured by LLaCTR at the field-level interactions, we analyzed the field heatmap of LLaCTR on the Video Games and MovieLens-1M dataset, as shown in Figure B.3.
-Figure B.3: Case study on the heatmap of LLM’s field embeddings.
-We can observe several interesting results: 1) For the Video Games dataset, there is a significant heat value between the field embeddings of "average_rating" and "rating_number", which aligns with the common sense that "users tend to purchase items with both high average ratings and high sales numbers".
-Increasing the interaction weight between these fields can help the CTR model make better predictions.
-Additionally, "average_rating" and "rating_number" exhibit high heat values with several other fields such as "item_id" and "user_id", indicating the relative importance of these fields.
-This insight could potentially inspire research in feature selection.
-2) For the MovieLens-1M dataset, the heat value between "UserID" and "MovieID" is relatively high, which to some extent explains the significant individual preference differences in MovieLens-1M.
-Injecting this information to guide the learning of feature interactions can assist the model in better capturing the semantic information in the dataset.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The main claims made in the abstract accurately conclude this paper’s contribution, including the findings and experimental results.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: The limitations of this work is discussed in Section 6.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper does not include theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The paper fully discloses all the information needed to reproduce the experimental results in Section 4 and Appendix A.5.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: The data and code are provided in https://anonymous.4open.science/r/LLaCTR-EC46.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: All the experimental details are included in Section 4 and Appendix A.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: The $p$-value of the main performance is reported in Table 1.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: Sufficient information on the computer resources is provided in Figure 4 and Appendix B.1.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The research conducted in the paper conform in every respect with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: This work proposed an effective and efficient LLM-enhanced CTR method and has no societal impact.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This paper poses no risks for misuse.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: The original paper that produced the codes and datasets are cited in the paper without any omissions.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: The new assets introduced in the paper are well documented and the details of the code are included.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: The LLM is used only for writing and editing in this paper.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+(a) $\quad\quad\lambda_{kl}$
+(b) $\quad\quad\lambda_{kl}$
+(c) $\quad\quad\lambda_{kl}$
+(d) $\quad\quad\lambda_{fm}$
+(e) $\quad\quad\lambda_{fm}$
+(f) $\quad\quad\lambda_{fm}$
+Figure 6. Sensitivity analysis w.r.t. $\lambda_{kl}$, $\lambda_{fm}$.